### v2

# Stacked Pseudo-Convergent Sequences and Polynomial Dedekind Domains

Abstract
Let $p\in\mathbb{Z}$ be a prime, $\overline{\mathbb{Q}_{p}}$ a fixed algebraic closure of the field of $p$-adic numbers and $\overline{\mathbb{Z}_{p}}$ the absolute integral closure of the ring of $p$-adic integers. Given a residually algebraic torsion extension $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, by Kaplansky’s characterization of immediate extensions of valued fields, there exists a pseudo-convergent sequence of transcendental type $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{(p),E}=\{\phi\in\mathbb{Q}(X)\mid\phi(s_{n})\in\overline{\mathbb{Z}_{p}},\text{ for all sufficiently large }n\in\mathbb{N}\}$. We show here that we may assume that $E$ is stacked, in the sense that, for each $n\in\mathbb{N}$, the residue field (the value group, respectively) of $\overline{\mathbb{Z}_{p}}\cap\mathbb{Q}_{p}(s_{n})$ is contained in the residue field (the value group, respectively) of $\overline{\mathbb{Z}_{p}}\cap\mathbb{Q}_{p}(s_{n+1})$; this property of $E$ allows us to describe the residue field and value group of $W$. In particular, if $W$ is a DVR, then there exists $\alpha$ in the completion $\mathbb{C}_{p}$ of $\overline{\mathbb{Q}_{p}}$, $\alpha$ transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$, where $\mathbb{O}_{p}$ is the unique local ring of $\mathbb{C}_{p}$; $\alpha$ belongs to $\overline{\mathbb{Q}_{p}}$ if and only if the residue field extension $W/M\supseteq\mathbb{Z}/p\mathbb{Z}$ is finite. As an application, we provide a full characterization of the Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$.
Keywords: pseudo-convergent sequence, residually algebraic extension, distinguished pair, minimal pair, Dedekind domain.
MSC Primary 12J20, 13F30, 13A18, 13F05, 13B25, 13F20.
Introduction
The problem of characterizing the set of the extensions of a valuation domain $V$ with quotient field $K$ to the field of rational functions $K(X)$ has a long and rich tradition (for example, see [1, 2, 3, 4, 13, 16, 21, 22, 23]). One of the recent direction of research is to describe these extensions by means of pseudo-monotone sequences of $K$ ([23]), in the original spirit of Ostrowski (see [18]) who introduced the well-known notion of pseudo-convergent sequence, later on expanded by Kaplansky in [13] to study immediate extensions of valued fields.
Here, given a prime $p\in\mathbb{Z}$ and the DVR $\mathbb{Z}_{(p)}$ of $\mathbb{Q}$, we are interested in describing residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, that is, valuation domains $W$ of $\mathbb{Q}(X)$ lying above $\mathbb{Z}_{(p)}$ such that the residue field extension $W/M\supseteq\mathbb{Z}/p\mathbb{Z}$ is algebraic and the value group $\Gamma_{w}$ of the associated valuation $w$ to $W$ is contained in the divisible hull of the value group of $\mathbb{Z}_{(p)}$ (i.e., the rationals). These valuation domains arise naturally as overrings of rings of integer-valued polynomials and Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ (see [9, 19]) and also in the description of closed subfields of $\mathbb{C}_{p}$ ([12]), the completion of an algebraic closure $\overline{\mathbb{Q}_{p}}$ of the field of $p$-adic numbers $\mathbb{Q}_{p}$. In the case $W$ is a DVR and the residue field extension is finite, then by [21, Theorem 2.5 & Proposition 2.2] there exists an element $\alpha$ in $\overline{\mathbb{Q}_{p}}$, transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\overline{\mathbb{Z}_{p}}\}$, where $\overline{\mathbb{Z}_{p}}$ is the absolute integral closure of $\mathbb{Z}_{p}$ (i.e., the integral closure of $\mathbb{Z}_{p}$ in $\overline{\mathbb{Q}_{p}}$; note that $\overline{\mathbb{Z}_{p}}$ is the valuation domain of the unique extension of $v_{p}$ to $\overline{\mathbb{Q}_{p}}$). In general, given a residually algebraic torsion extension $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, there exists a pseudo-convergent sequence $E=\{s_{n}\}_{n\in\mathbb{N}}$ in $\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{(p),E}=\{\phi\in\mathbb{Q}(X)\mid\phi(s_{n})\in\overline{\mathbb{Z}_{p}},\text{ for all sufficiently large }n\in\mathbb{N}\}$ (Proposition 2.24). One of the main result of this paper is to show that we may assume that $E$ is stacked (in a sense we make clear in §2; see Theorem 2.5). In particular, if $W$ is a DVR of $\mathbb{Q}(X)$ extending $\mathbb{Z}_{(p)}$ such that the extension of the residue fields is infinite algebraic, then there exists $\alpha$ in $\mathbb{C}_{p}\setminus\overline{\mathbb{Q}_{p}}$, such that $W=\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$, where $\mathbb{O}_{p}$ is the completion of $\overline{\mathbb{Z}_{p}}$ (equivalently, $\mathbb{O}_{p}$ is the valuation domain of the unique extension of $v_{p}$ to $\mathbb{C}_{p}$; see Corollary 2.28). Necessarily, the (transcendental) extension $\mathbb{Q}_{p}(\alpha)/\mathbb{Q}_{p}$ has finite ramification.
It is worth recalling that in [4, §5, 1., & Theorem 5.1] a residually algebraic torsion extension $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ is realized as the limit of a sequence of residually transcendental extensions $W_{n}$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ (i.e., the residue field extension of $W_{n}$ over $\mathbb{Z}_{(p)}$ is transcendental); moreover, for each $n\in\mathbb{N}$, $W_{n}$ is defined by a minimal pair $(s_{n},\delta_{n})$ (as explained in [4, p. 282]; for the definition of minimal pair see §1.2). Here, $W$ is realized as the valuation domain $\mathbb{Z}_{(p),E}$, where for each $n\in\mathbb{N}$, $(s_{n},\delta_{n}=v_{p}(s_{n+1}-s_{n}))$ is a minimal pair, too.
The motivations behind these results are based on the paper [5], in which the authors study closed subfields of $\mathbb{C}_{p}$ and show that any transcendental element of $\mathbb{C}_{p}$ is the limit of a particular kind of Cauchy sequence in $\overline{\mathbb{Q}_{p}}$ called distinguished ([5, Proposition 2.2]) which allows them to associate to such an element a set of invariants ([5, Remark 2.4]). The notion of stacked sequence we introduce in this paper is a generalization of the notion of distinguished sequence and falls into the well-known class of pseudo-convergent sequences. It allows us to describe the whole class of residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, which strictly comprise the valuation domains $\mathbb{Z}_{(p),\alpha}$ arising from elements $\alpha\in\mathbb{C}_{p}\setminus\overline{\mathbb{Q}_{p}}$.
As an application of the above results, we are able to complete the classification of the family of Dedekind domains $R$ between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ started in [19]. In that paper we described the Dedekind domains of this family whose residue fields of prime characteristic are finite fields ([19, Theorem 2.17]); the description is obtained by means of the notion of rings of integer-valued polynomials over algebras. We also showed that, given a group $G$ which is the direct sum of a countable family of finitely generated abelian groups, there exists a Dedekind domain $R$ with finite residue fields of prime characteristic, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, with class group $G$ ([19, Theorem 3.1]).
The paper is organized as follows. In Section 1 we recall the relevant notions we need in our paper: first, we review the definition of pseudo-convergent sequence of a valued field $K$ and the valuation domain of $K(X)$ associated to such a sequence in the spirit of Ostrowski ([18]), as developed recently in [22, 23]. Then, we recall the notion of distinguished pair introduced in [25] which later on was used in [5] to describe closed subfield of $\mathbb{C}_{p}$ in terms of a specific kind of pseudo-convergent Cauchy sequence called distinguished.
In Section 2, we introduce the notion of stacked sequence $E=\{s_{n}\}_{n\in\mathbb{N}}$ in $\overline{\mathbb{Q}_{p}}$, which turns out to be a pseudo-convergent sequence of transcendental type such that, for each $n\in\mathbb{N}$, the value group (the residue field, respectively) of $\overline{\mathbb{Z}_{p}}\cap\mathbb{Q}_{p}(s_{n})$ is contained in the value group (the residue field, respectively) of $\overline{\mathbb{Z}_{p}}\cap\mathbb{Q}_{p}(s_{n+1})$. By Theorem 2.5, every residually algebraic extension $W$ of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ can be realized by means of a stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$, that is, $W=\mathbb{Z}_{p,E}=\{\phi\in\mathbb{Q}_{p}(X)\mid\phi(s_{n})\in\overline{\mathbb{Z}_{p}},\text{ for all sufficiently large }n\in\mathbb{N}\}$. Moreover, the above specific property of stacked sequences is crucial for the description of the residue field and value group of $W$ as the union of the ascending chain of residue fields and value groups of $\overline{\mathbb{Z}_{p}}\cap\mathbb{Q}_{p}(s_{n})$, respectively (Proposition 2.7). We mentioned above that the elements $\alpha\in\mathbb{C}_{p}\setminus\overline{\mathbb{Q}_{p}}$ such that the extension $\mathbb{Q}_{p}(\alpha)/\mathbb{Q}_{p}$ has finite ramification give raise to DVRs of $\mathbb{Q}(X)$; we characterize such elements as the limits of sequences contained in the maximal unramified extension of a finite extension of $\mathbb{Q}_{p}$ (Proposition 2.20). We close this section by pointing out a wrong statement in the paper [12], namely, that the completion of $\mathbb{Q}_{p}(X)$ with respect to a residually algebraic torsion extension $W$ of $\mathbb{Z}_{p}$ is a subfield of $\mathbb{C}_{p}$; this is not true in general and it depends on whether the above sequence $E$ is Cauchy or not. In §2.3, we use the result of §2.1 about residually algebraic torsion extensions of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ in order to characterize the analogous extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ (Proposition 2.24). In Theorem 2.26, we show that, for any prescribed algebraic extension $k$ of $\mathbb{F}_{p}$ and value group $\Gamma$, $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$, there exists $\alpha\in\mathbb{C}_{p}$, transcendental over $\mathbb{Q}$, such that $\mathbb{Z}_{(p),\alpha}$ has residue field $k$ and value group $\Gamma$.
Finally, in Section 3 we provide the aforementioned classification of the Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ by means of the notion of ring of integer-valued polynomials over an algebra: given such a domain $R$, we show that for each prime $p\in\mathbb{Z}$ there exists a finite set $E_{p}\subset\mathbb{C}_{p}$ of transcendental elements over $\mathbb{Q}$ such that $R=\{f\in\mathbb{Q}[X]\mid f(E_{p})\subseteq\mathbb{O}_{p},\forall p\in\mathbb{P}\}$ (Theorem 3.4).
1 Preliminaries
We refer to [6, 10, 26, 27] for generalities about valuation theory. A valuation domain $W$ of the field of rational functions $K(X)$ is an extension of a valuation domain $V$ of $K$ if $W\cap K=V$. We denote by $w$ a valuation associated to $W$, by $\Gamma_{w}$ the value group of $w$ and by $k_{w}$ the residue field of $W$. We recall that an extension $W$ of $V$ to $K(X)$ is called residually algebraic if the residue field extension is algebraic and it is called torsion if $\Gamma_{w}$ is contained in the divisible hull of the value group $\Gamma_{v}$ of $V$ (see [4]). Given a valuation domain $W$ with quotient field $F$, a subfield $K$ of $F$ and the valuation domain $V=W\cap K$, we say that $W$ is an immediate extension of $V$ (or simply immediate over $V$) if the value groups (the residue fields, respectively) of $V$ and $W$ are the same. Given a field $K$ with a valuation domain $V$, we denote by $\widehat{K}$ ($\widehat{V}$, respectively) the completion of $K$ ($V$, respectively) with respect to $V$-adic topology.
1.1 Pseudo-convergent sequences
The following basic material about pseudo-convergent sequences can be found for example in [13, 22, 23].
Given a valued field $(K,v)$, a sequence $E=\{s_{n}\}_{n\in\mathbb{N}}\subset K$ is said to be pseudo-convergent if, for all $n<m<k$ we have
$v(s_{n}-s_{m})<v(s_{m}-s_{k}).$
In particular, for all $n$ and $m>n$ we have $v(s_{n}-s_{m})=v(s_{n}-s_{n+1})$. For each $n\in\mathbb{N}$, we set $\delta_{n}=v(s_{n}-s_{n+1})$. The strictly increasing sequence $\{\delta_{n}\}_{n\in\mathbb{N}}$ of the value group $\Gamma_{v}$ of $v$ is called the gauge of $E$. The sequence $E$ is a classical Cauchy sequence in $K$ if and only if the gauge of $E$ is cofinal in $\Gamma_{v}$. In this case, $E$ converges to a unique limit $\alpha\in\widehat{K}$. In general, if $E=\{s_{n}\}_{n\in\mathbb{N}}\subset K$ is a pseudo-convergent sequence, we say that an element $\alpha\in K$ is a pseudo-limit of $E$ if $v(s_{n}-\alpha)$ is a strictly increasing sequence. Equivalently, $v(s_{n}-\alpha)=\delta_{n}$ for each $n\in\mathbb{N}$. The set of pseudo-limits $\mathcal{L}_{E}$ in $K$ of a pseudo-convergent sequence $E$ is equal to $\mathcal{L}_{E}=\alpha+\text{Br}(E)$ ([13, Lemma 3]), where $\text{Br}(E)=\{x\in K\mid v(x)>\delta_{n},\forall n\in\mathbb{N}\}$ is a fractional ideal, called the breadth ideal of $E$. Clearly, $E$ is a Cauchy sequence if and only if $\text{Br}(E)=\{0\}$.
As in [13, Definitions, p. 306], a pseudo-convergent sequence $E=\{s_{n}\}_{n\in\mathbb{N}}\subset K$ is of transcendental type if, for all $f\in K[X]$, $v(f(s_{n}))$ is eventually constant. Otherwise, $E$ is said to be of algebraic type if $v(f(s_{n}))$ is eventually strictly increasing for some $f\in K[X]$. The sequence $E$ is of algebraic type if and only if, for some extension $u$ of $v$ to the algebraic closure $\overline{K}$ of $K$, there exists $\alpha\in\overline{K}$ which is a pseudo-limit of $E$ with respect to $u$. If $F$ is a subfield of $K$, then we say that $E$ is of transcendental type over $F$ if, for all $f\in F[X]$, $v(f(s_{n}))$ is eventually constant. Almost all the pseudo-convergent sequences considered in this paper in order to describe residually algebraic torsion extensions to the field of rational functions are of transcendental type.
Given a pseudo-convergent sequence $E=\{s_{n}\}_{n\in\mathbb{N}}\subset K$, the following is a valuation domain of $K(X)$ extending $V$ associated to $E$ ([22, Theorem 3.8]):
$V_{E}=\{\phi\in K(X)\mid\phi(s_{n})\in V,\text{ for all sufficiently large }n\in\mathbb{N}\}.$
Moreover, by the same Theorem, $X$ is a pseudo-limit of $E$ with respect to the valuation $v_{E}$ associated to $V_{E}$, so, in particular, $v_{E}(X-s_{n})=\delta_{n}$, for every $n\in\mathbb{N}$. Also, if $E$ is of transcendental type, then for all $f\in K[X]$, we have $v_{E}(f)=v(f(s_{n}))$ for all $n$ sufficiently large ([13, Theorem 2] or [22, Theorem 4.9, a)]).
In case $E$ is a Cauchy sequence converging to $\alpha\in\widehat{K}$, then
$V_{E}=V_{\alpha}=\{\phi\in K(X)\mid\phi(\alpha)\in\widehat{V}\}$
(see [22, Remark 3.10]).
Given two pseudo-convergent sequences $E=\{s_{n}\}_{n\in\mathbb{N}},E^{\prime}=\{s_{n}^{\prime}\}_{n\in\mathbb{N}}\subset K$, we say that $E,E^{\prime}$ are equivalent if $\text{Br}(E)=\text{Br}(E^{\prime})$ and for each $k\in\mathbb{N}$, there exist $i_{0},j_{0}\in\mathbb{N}$ such that $v(s_{i}-s_{j}^{\prime})>v(s_{k+1}-s_{k})$ for each $i\geq i_{0}$ and $j\geq j_{0}$ (see [22, §5]). By [23, Proposition 5.3], $E,E^{\prime}$ are equivalent if and only if $V_{E}=V_{E^{\prime}}$.
1.2 Distinguished pairs
We suppose in this section that $(K,v)$ is a complete valued field where $v$ is a rank one discrete valuation (so, in particular, $(K,v)$ is Henselian). Let $\overline{K}$ be a fixed algebraic closure and let $v$ denote the unique extension of $v$ to $\overline{K}$. Let also $\Gamma_{\overline{v}}=\Gamma_{v}\otimes\mathbb{Q}$ be the divisible hull of $\Gamma_{v}$. Given an element $a\in\overline{K}$, let $O_{a},k_{a},\Gamma_{a}$ be the valuation domain of the restriction of $v$ to $K(a)$, the residue field of $O_{a}$ and the value group of $O_{a}$, respectively.
As in [14], given $a\in\overline{K}\setminus K$, we set:
$\delta_{K}(a)=$ $\sup\{v(a-c)\mid c\in\overline{K},[K(c):K]<[K(a):K]\}$
$\omega_{K}(a)=$ $\sup\{v(a-a^{\prime})\mid a^{\prime}\not=a\text{ runs over the }K\text{-conjugates of }a\}$
The following is the well-known Krasner’s lemma. Essentially, given a separable element $a\in\overline{K}$, if another element $b\in\overline{K}$ is closer to $a$ than to any other of its conjugates, then $K(a)$ is a subfield of $K(b)$.
Lemma 1.1 (Krasner).
If $a\in\overline{K}^{\text{sep}}$ and $b\in\overline{K}$ is such that $v(a-b)>\omega_{K}(a)$, then $K(a)\subseteq K(b)$.
In particular, for every $a\in\overline{K}^{\text{sep}}$ we have $\delta_{K}(a)\leq\omega_{K}(a)$. Moreover, it follows also that $\delta_{K}(a)$ is a maximum, since $v$ is supposed to be discrete. This is known (see for example [25, p. 105]), but for the sake of the reader we give a short proof.
Lemma 1.2.
In the above setting, $\delta_{K}(a)=\max\{v(a-c)\mid c\in\overline{K},[K(c):K]<[K(a):K]\}$.
Proof.
By Krasner’s Lemma, for each of the relevant $c$’s we have $v(a-c)\leq\omega_{K}(a)$. Note that the ramification index of $K(a,c)$ over $K$ is (strictly) bounded by $[K(a):K]^{2}$. In particular, since the value $v(a-c)$ belongs to $\Gamma_{a-c}$, it follows that there exists $N\in\mathbb{N}$, independent from each of the above $c$’s, such that $Nv(a-c)\in\Gamma_{v}\cong\mathbb{Z}$. Hence, the set $\{v(a-c)\mid c\in\overline{K},[K(c):K]<[K(a):K]\}$ (which is a subset of $\Gamma_{\overline{v}}$) is bounded from above and its elements have bounded torsion. It follows that this set has a maximum, which is equal to $\delta_{K}(a)$ by its very definition. ∎
Similar to Krasner’s lemma, we have the following fundamental principle (see [14, Theorem 1.1]), first discovered in [25].
Theorem 1.3.
Suppose that $a,b\in\overline{K}$ are such that $v(a-b)>\delta_{K}(b)$. Then:
  • i)
    $\Gamma_{b}\subseteq\Gamma_{a}$.
  • ii)
    $k_{b}\subseteq k_{a}$.
  • iii)
    $[K(b):K]\mid[K(a):K]$.
Next, we recall the definition of distinguished pair introduced in [25, p. 105].
Definition 1.4.
A pair of elements $(b,a)\in\overline{K}^{2}$ is said to be distinguished if the following hold:
  • i)
    $[K(b):K]<[K(a):K]$.
  • ii)
[…]
  • iii)
[…]
Part of the definition of distinguished pair is related to the notion of minimal pair, which we now recall (see for example [4, 2, 3]).
Definition 1.5.
[…]
In other words, $(a,\delta)$ is a minimal pair if for every $b\in B(a,\delta)=\{x\in\overline{K}\mid v(a-x)\geq\delta\}$, we have $[K(b):K]\geq[K(a):K]$ (i.e., $a$ is a ’center’ of the ball $B(a,\delta)$ of minimal degree). By Lemma 1.2, $(a,\delta)$ is a minimal pair if and only if $\delta>\delta_{K}(a)$. In particular, if $\delta>\omega_{K}(a)$, then $(a,\delta)$ is a minimal pair.
Remarks 1.6.
[…]
Note that iii) is equivalent to the following:
  • iii’)
[…]
$v(a-b)>\delta_{K}(b).$
In particular, by the above Theorem, $\Gamma_{b}\subseteq\Gamma_{a}$, $k_{b}\subseteq k_{a}$ and $[K(b):K]\mid[K(a):K]$.
Finally, note also that $\delta_{K}(b)<\delta_{K}(a)$.
2 Stacked pseudo-convergent sequences of $\overline{\mathbb{Q}_{p}}$
Let $\mathbb{P}\subset\mathbb{Z}$ be the set of prime numbers and let $p\in\mathbb{P}$ be a fixed prime. We let $\mathbb{Z}_{(p)}$ be the localization of $\mathbb{Z}$ at the prime ideal $p\mathbb{Z}$, $\mathbb{Z}_{p}$ the ring of $p$-adic integers and $\mathbb{Q}_{p}$ its field of fractions, the field of $p$-adic numbers. If $v_{p}$ denotes the usual $p$-adic valuation, then $\mathbb{Z}_{p}$ ($\mathbb{Q}_{p}$, respectively) is the completion of $\mathbb{Z}$ ($\mathbb{Q}$, respectively) with respect to the $p$-adic valuation. We denote by $\overline{\mathbb{Q}_{p}}$ a fixed algebraic closure of $\mathbb{Q}_{p}$ and still denote the unique extension of $v_{p}$ to $\overline{\mathbb{Q}_{p}}$ by $v_{p}$. Note that $\overline{\mathbb{Q}_{p}}$ is a rank one non-discrete valued field with valuation domain denoted by $\overline{\mathbb{Z}_{p}}$, the integral closure of $\mathbb{Z}_{p}$ in $\overline{\mathbb{Q}_{p}}$. We will use the well-known fact that $\mathbb{Q}_{p}$ has only finitely many extensions of a given degree (see for example [17, Corollary 2, Chapter V, p. 202]).
Finally, we let $\mathbb{C}_{p}$ be the completion of $\overline{\mathbb{Q}_{p}}$ with respect to the $p$-adic valuation and by $\mathbb{O}_{p}$ the completion of $\overline{\mathbb{Z}_{p}}$; $v_{p}$ still denotes the unique extension of $v_{p}$ to $\mathbb{C}_{p}$. For $\alpha\in\overline{\mathbb{Q}_{p}}\setminus\mathbb{Q}_{p}$, for short we set $\delta_{\mathbb{Q}_{p}}(\alpha)=\delta(\alpha)$ and $\omega_{\mathbb{Q}_{p}}(\alpha)=\omega(\alpha)$. For $\alpha\in\mathbb{C}_{p}$, we denote by $e_{\alpha}$ ($f_{\alpha}$, respectively) the ramification index (the residue field degree, respectively) of $\mathbb{Q}_{p}(\alpha)$ over $\mathbb{Q}_{p}$. Clearly, if $\alpha\in\overline{\mathbb{Q}_{p}}$ then $e_{\alpha}\cdot f_{\alpha}<\infty$; we show that the converse holds in Remark 2.15. Note that each element of $\mathbb{C}_{p}\setminus\overline{\mathbb{Q}_{p}}$ is transcendental over $\mathbb{Q}_{p}$; we call such elements simply transcendental. For a transcendental element $\alpha\in\mathbb{C}_{p}$, even if $e_{\alpha}\cdot f_{\alpha}=\infty$, we will show in Theorem 2.21 that either one of $e_{\alpha}$ or $f_{\alpha}$ can be finite.
2.1 Residually algebraic torsion extensions of $\mathbb{Z}_{p}$
In this section we describe residually algebraic torsion extensions of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ by means of a suitable class of pseudo-convergent sequences of transcendental type contained in $\overline{\mathbb{Q}_{p}}$, called stacked sequence, which we now introduce. This definition is a generalization of [5, p. 135] [^2] The notion of distinguished sequence was introduced in [5]. We cannot borrow that term here for our sequences for the following reason: by Lemma 2.3, a stacked sequence is pseudo-convergent, and distinguished pseudo-convergent sequences have already been defined by P. Ribenboim on p. 474 of Corps maximaux et complets par des valuations de Krull. Math. Z. 69 (1958), 466–479, to denote pseudo-convergent sequences of a valued field whose breadth ideal is a non-maximal prime ideal..
Definition 2.1.
Let $E=\{s_{n}\}_{n\geq 0}\subset\overline{\mathbb{Q}_{p}}$ be a sequence with $s_{0}\in\mathbb{Q}_{p}$. For every $n\geq 0$, we consider the following properties:
  • i)
    $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$.
  • ii)
    for every $c\in\overline{\mathbb{Q}_{p}}$ such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$, $v(s_{n+1}-c)\leq v(s_{n+1}-s_{n})$.
  • iii)
    for every $c\in\overline{\mathbb{Q}_{p}}$ such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$, $v(s_{n}-c)<v(s_{n+1}-s_{n})$.
We say that $E$ is unbounded if i) holds for every $n$, stacked if i) and iii) hold for every $n$, and strongly stacked if i), ii), iii) hold for every $n$. Equivalently, $E$ is stacked if i) holds and $(s_{n},\delta_{n}=v(s_{n+1}-s_{n}))$ is a minimal pair for every $n\geq 0$ and $E$ is strongly stacked if $(s_{n},s_{n+1})$ is distinguished for every $n\geq 0$.
Remark 2.2.
Let $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ be a stacked sequence. Note that the sequence $\{v(s_{n+1}-s_{n})=\delta_{n}\}_{n\in\mathbb{N}}$ is strictly increasing, since $[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and $(s_{n},\delta_{n})$ is a minimal pair. In the original definition of distiguished sequence $E$ in [5], the sequence $\{\delta_{n}\}_{n\in\mathbb{N}}$ is unbounded, thus, in this case $E$ is a Cauchy sequence. In our setting we are not imposing that restriction; we show in Lemma 2.3 below that a stacked sequence is a pseudo-convergent sequence of transcendental type of $\overline{\mathbb{Q}_{p}}$.
The motivation for the terminology of these kind of sequences, is due to the following fact. For each $n\in\mathbb{N}$, for short we set $\Gamma_{n}=\Gamma_{s_{n}}$ and $k_{n}=k_{s_{n}}$ (i.e., the value group and the residue field of the valuation domain $O_{s_{n}}$ of $\mathbb{Q}_{p}(s_{n})$, respectively). By Remarks 1.6, $v(s_{n+1}-s_{n})>\delta_{K}(s_{n})$. Hence, by Theorem 1.3, we have $\Gamma_{n}\subseteq\Gamma_{n+1}$ and $k_{n}\subseteq k_{n+1}$. For each $n\in\mathbb{N}$, we set $e_{n}=e(\mathbb{Q}_{p}(s_{n})\mid\mathbb{Q}_{p})$ and $f_{n}=f(\mathbb{Q}_{p}(s_{n})\mid\mathbb{Q}_{p})$, the ramification index and the residue field degree of $O_{s_{n}}$ over $\mathbb{Z}_{p}$, respectively; we remark that $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=e_{n}f_{n}=d_{n}$ for each $n\in\mathbb{N}$, and since $\{d_{n}\}_{n\in\mathbb{N}}$ is unbounded by assumption, either $\{e_{n}\}_{n\in\mathbb{N}}$ is unbounded or $\{f_{n}\}_{n\in\mathbb{N}}$ is unbounded. Since $e_{n}\mid e_{n+1}$ for each $n\in\mathbb{N}$, $\{e_{n}\}_{n\in\mathbb{N}}$ is bounded if and only if $e_{n}=e$ for all $n\in\mathbb{N}$ sufficiently large. Similarly for $\{f_{n}\}_{n\in\mathbb{N}}$.
By Remarks 1.6, condition ii) is equivalent to $\delta_{n}=v(s_{n}-s_{n+1})=\delta_{K}(s_{n+1})$ (note that in general the inequality $\delta_{n}\leq\delta_{K}(s_{n+1})$ holds). In other words, among all the elements $c\in\overline{\mathbb{Q}_{p}}$ such that $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]\leq[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$, $s_{n}$ is one of those which is closest to $s_{n+1}$.
Let $E=\{t_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ be a pseudo-convergent sequence. If $\{[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]\mid n\in\mathbb{N}\}$ is bounded, then $E$ is contained in a finite extension $K$ of $\mathbb{Q}_{p}$ and hence $E$ is Cauchy and therefore converges to an element $\alpha\in K$. In particular, if $E$ is of transcendental type, then the set $\{[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]\}_{n\in\mathbb{N}}$ is necessarily unbounded. Stacked sequences are of this kind, as the next lemma shows.
Lemma 2.3.
Let $E\subset\overline{\mathbb{Q}_{p}}$ be a stacked sequence. Then $E$ is a pseudo-convergent sequence of transcendental type.
Proof.
Let $E=\{s_{n}\}_{n\in\mathbb{N}}$ and set $\delta_{n}=v(s_{n+1}-s_{n})$, for each $n\in\mathbb{N}$. We have already observed in Remark 2.2 that $\{\delta_{n}\}_{n\in\mathbb{N}}$ is a strictly increasing sequence. Moreover, for every $m>n$, we have $v(s_{n}-s_{m})>v(s_{n}-s_{n-1})$. In particular, $v(s_{n-1}-s_{m})=v(s_{n-1}-s_{n})$ for every $m\geq n$. Let now $n<m<k$. Then
$v(s_{n}-s_{m})=v(s_{n}-s_{n+1})<v(s_{m}-s_{m+1})=v(s_{m}-s_{k})$
which shows that $E$ is a pseudo-convergent sequence.
We prove now that $E$ is of transcendental type. Let $\alpha\in\overline{\mathbb{Q}_{p}}$. Then there exists $n\in\mathbb{N}$ such that $[\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Since $(s_{n},\delta_{n})$ is a minimal pair, $v(s_{n}-\alpha)<\delta_{n}$, so, in particular, $\alpha$ cannot be a pseudo-limit of $E$. This shows that $E$ has no pseudo-limits in $\overline{\mathbb{Q}_{p}}$, thus $E$ is of transcendental type. ∎
Let $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ be a stacked sequence. In particular, by Lemma 2.3, the sequence $\{\delta_{n}=v(s_{n+1}-s_{n})\}_{n\in\mathbb{N}}$ is the gauge of the pseudo-convergent sequence $E$. Moreover, by the same Lemma, if $E$ is Cauchy, then $E$ converges to a transcendental element $\alpha\in\mathbb{C}_{p}$.
The next proposition shows that any residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ is obtained by means of a pseudo-convergent sequence of transcendental type of $\overline{\mathbb{Q}_{p}}$. We recall that if $E\subset\overline{\mathbb{Q}_{p}}$ is a pseudo-convergent sequence of transcendental type, then $\overline{\mathbb{Z}_{p}}_{E}$, the associated valuation domain of $\overline{\mathbb{Q}_{p}}(X)$, is an immediate extension of $\overline{\mathbb{Z}_{p}}$ and conversely every immediate extension of $\overline{\mathbb{Z}_{p}}$ to $\overline{\mathbb{Q}_{p}}(X)$ can be realized in this way (see for example [13, 23]). If $\mathbb{Z}_{p,E}=\overline{\mathbb{Z}_{p}}_{E}\cap\mathbb{Q}_{p}(X)$, then $\mathbb{Z}_{p,E}$ is a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$.
Proposition 2.4.
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$. Then there exists a pseudo-convergent sequence $E\subset\overline{\mathbb{Q}_{p}}$ of transcendental type such that
$W=\mathbb{Z}_{p,E}=\{\phi\in\mathbb{Q}_{p}(X)\mid\phi(s_{n})\in\overline{\mathbb{Z}_{p}},\text{ for all sufficiently large }n\in\mathbb{N}\}.$
Proof.
Let $\overline{W}$ be an extension of $W$ to $\overline{\mathbb{Q}_{p}}(X)$. Then $\overline{W}$ is an immediate extension of $\overline{\mathbb{Z}_{p}}$ to $\overline{\mathbb{Q}_{p}}(X)$ (and, in particular, is a residually algebraic torsion extension of $\overline{\mathbb{Z}_{p}}$). By [13, Theorems 1 & 2] or [23, Theorem 6.2, a)], there exists a pseudo-convergent sequence $E\subset\overline{\mathbb{Q}_{p}}$ of transcendental type such that $\overline{W}=\overline{\mathbb{Z}_{p}}_{E}$. The claim follows contracting down to $\mathbb{Q}_{p}(X)$. ∎
Clearly, not every pseudo-convergent sequence of transcendental type in $\overline{\mathbb{Q}_{p}}$ is stacked. However, the next theorem is the converse of Lemma 2.3: it shows that any pseudo-convergent sequence of transcendental type is equivalent to a strongly stacked sequence. In particular, every stacked sequence is equivalent to a strongly stacked sequence. Moreover, given a valuation domain $\mathbb{Z}_{p,E}$ of $\mathbb{Q}_{p}(X)$ associated to a pseudo-convergent sequence $E\subset\overline{\mathbb{Q}_{p}}$ of transcendental type, without loss of generality we may also assume that $E$ is strongly stacked.
By [5, Proposition 2.2], every transcendental element $t\in\mathbb{C}_{p}$ is the limit of a strongly stacked sequence $E$ of $\overline{\mathbb{Q}_{p}}$. The next theorem is the analogous of that result for residually algebraic extensions $W$ of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$: for such a valuation $W$, there exists a strongly stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{p,E}$; it is not difficult to show that, for a transcendental element $t\in\mathbb{C}_{p}$, the valuation domain $\mathbb{Z}_{p,t}=\{\phi\in\mathbb{Q}_{p}(X)\mid\phi(t)\in\mathbb{O}_{p}\}$ is a residually algebraic torsion extension of $\mathbb{Z}_{p}$.
Theorem 2.5.
Let $E\subset\overline{\mathbb{Q}_{p}}$ be a pseudo-convergent sequence of transcendental type. Then there exists a strongly stacked sequence $E^{\prime}\subset\overline{\mathbb{Q}_{p}}$ which is equivalent to $E$. In particular, given a residually algebraic torsion extension $W$ of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$, there exists a strongly stacked sequence $E^{\prime}\subset\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{p,E^{\prime}}$.
Proof.
Let $E=\{t_{n}\}_{n\in\mathbb{N}}$ and let $\overline{v}_{E}$ be a valuation associated to $\overline{\mathbb{Z}_{p}}_{E}\subset\overline{\mathbb{Q}_{p}}(X)$.
First, we consider the following subset of $\Gamma_{v_{E}}\subseteq\mathbb{Q}$:
$M_{E}(X,\mathbb{Q}_{p})=\{v_{E}(X-s)\mid s\in\mathbb{Q}_{p}\}.$
If $M_{E}(X,\mathbb{Q}_{p})$ is not bounded, then there exists a sequence $\{s_{n}\}_{n\in\mathbb{N}}\subset\mathbb{Q}_{p}$ such that $v_{E}(X-s_{n})$ tends to $\infty$. Necessarily, the sequence $\{s_{n}\}_{n\in\mathbb{N}}$ is Cauchy and so converges to an element $s$ of $\mathbb{Q}_{p}$. Now, for every $n$, $v_{E}(X-s_{n})=\overline{v}_{E}(X-s_{n})=v(t_{m}-s_{n})$ for all $m$ sufficiently large since $E$ is of transcendental type (see §1.1). Hence, $E$ would be a Cauchy sequence equivalent to $\{s_{n}\}_{n\in\mathbb{N}}$ and $E$ would converge to $s$, too, which is not possible. Let then $\delta_{0}=\sup M_{E}(X,\mathbb{Q}_{p})\in\mathbb{R}$. We claim that $\delta_{0}\in M_{E}(X,\mathbb{Q}_{p})$, that is, $\delta_{0}$ is a maximum. Suppose otherwise: there exists a sequence $\{r_{k}\}_{k\in\mathbb{N}}\subset\mathbb{Q}_{p}$ such that $v_{E}(X-r_{k})\nearrow\delta_{0}$. Then $\{r_{k}\}_{k\in\mathbb{N}}\subset\mathbb{Q}_{p}$ would be a pseudo-convergent sequence which is not Cauchy, which is not possible since $\mathbb{Q}_{p}$ is a complete valued field. Hence, there exists $s_{0}\in\mathbb{Q}_{p}$ such that $v_{E}(X-s_{0})=\delta_{0}$.
For $n>0$, we now choose $s_{n}\in\overline{\mathbb{Q}_{p}}$ so that $(s_{n-1},s_{n})$ is distinguished. Let $B_{n}$ be the subset of $\alpha$’s in $\overline{\mathbb{Q}_{p}}$ satisfying the following properties:
  • i)
    $[\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$.
  • ii)
    $\overline{v}_{E}(X-\alpha)>\overline{v}_{E}(X-s_{n-1})$.
  • iii)
    the positive integer $[\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]-[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ is minimal.
Note that since $\mathbb{N}$ is well-ordered, condition iii) can be satisfied (that is, among the $\alpha\in\overline{\mathbb{Q}_{p}}$ satisfying i) and ii), we can find one which also satisfies iii) ). Since $\overline{v}_{E}(X-s_{n-1})=v(t_{m}-s_{n-1})$ for all $m$ sufficiently large, for such $m$’s we also have $\overline{v}_{E}(X-s_{n-1})<\overline{v}_{E}(X-t_{m})$. Moreover, without loss of generality, we may also assume that $[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ since $\{[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]\}_{m\in\mathbb{N}}$ is unbounded. This shows that the set $B_{n}$ is non-empty. Let $M_{E}(X,B_{n})=\{\overline{v}_{E}(X-\alpha)\mid\alpha\in B_{n}\}$, which is a subset of $\mathbb{Q}$. Let $\delta_{n}=\sup M_{E}(X,B_{n})$. Since each element of $B_{n}$ has the same degree over $\mathbb{Q}_{p}$, it follows that $B_{n}$ is contained in a finite extension $K$ of $\mathbb{Q}_{p}$. In particular, it follows as above that $M_{E}(X,B_{n})$ is upper bounded. Let $\delta_{n}=\sup M_{E}(X,B_{n})\in\mathbb{R}$. Next, we show that $M_{E}(X,B_{n})$ contains its upper bound (which is, in particular, a rational number). Suppose otherwise: then there exists a sequence $\{\alpha_{k}\}_{k\in\mathbb{N}}\subset B_{n}$ such that $\overline{v}_{E}(X-\alpha_{k})\nearrow\delta_{n}$. In particular, $\{\alpha_{k}\}_{k\in\mathbb{N}}$ would be a pseudo-convergent sequence of a finite extension of $\mathbb{Q}_{p}$ which is not Cauchy, which is impossible. Let $s_{n}\in B_{n}$ be such that $\overline{v}_{E}(X-s_{n})=\delta_{n}$. Note that $v_{p}(s_{n}-s_{n-1})=\overline{v}_{E}(s_{n}-X+X-s_{n-1})=\overline{v}_{E}(X-s_{n-1})=\delta_{n-1}$.
We now show that $(s_{n-1},s_{n})$ is distinguished. Clearly, $[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$.
Let $c\in\overline{\mathbb{Q}_{p}}$ be such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$.
If $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$, then by the minimality of the degree of $s_{n}$, we have $\overline{v}_{E}(X-c)\leq\overline{v}_{E}(X-s_{n-1})=\delta_{n-1}$, so
$v_{p}(s_{n}-c)=\overline{v}_{E}(s_{n}-X+X-c)=\overline{v}_{E}(X-c)\leq\delta_{n-1}=v_{p}(s_{n}-s_{n-1})$
Suppose now that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$.
If $\overline{v}_{E}(X-c)\leq\overline{v}_{E}(X-s_{n-2})$, then $\overline{v}_{E}(X-c)<\delta_{n-1}$.
If $\overline{v}_{E}(X-c)>\overline{v}_{E}(X-s_{n-2})$ then $c\in B_{n-1}$, so $\overline{v}_{E}(X-c)\leq\delta_{n-1}=\overline{v}_{E}(X-s_{n-1})$.
In either case,
$v_{p}(s_{n}-c)=\overline{v}_{E}(s_{n}-X+X-c)=\overline{v}_{E}(X-c)\leq\delta_{n-1}=v_{p}(s_{n}-s_{n-1}).$
Note that in particular, for $n=1$ we have that $(s_{0},s_{1})$ is distinguished, since condition iii) of Definition 2.1 is empty, since $s_{0}\in\mathbb{Q}_{p}$.
Suppose now that $n\geq 2$ and assume by induction that $(s_{n-2},s_{n-1})$ is distinguished. Let $c\in\overline{\mathbb{Q}_{p}}$ be such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$. Since $(s_{n-2},s_{n-1})$ is distinguished, we have $v_{p}(s_{n-1}-c)\leq v_{p}(s_{n-1}-s_{n-2})=\delta_{n-2}<\delta_{n-1}$. Hence,
$v_{p}(s_{n}-c)=v_{p}(s_{n}-s_{n-1}+s_{n-1}-c)=v_{p}(s_{n-1}-c)<v_{p}(s_{n}-s_{n-1}).$
We now show that $E^{\prime}=\{s_{n}\}_{n\in\mathbb{N}}$ is equivalent to $E=\{t_{n}\}_{n\in\mathbb{N}}$. Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ and $\{\delta_{n}\}_{n\in\mathbb{N}}$ be the gauges of $E$ and $E^{\prime}$, respectively. We need to show that, for each $k\in\mathbb{N}$, there exists $n\in\mathbb{N}$ such that $\lambda_{k}\leq\delta_{n}$. Since $E^{\prime}$ is unbounded, there exists $n\in\mathbb{N}$ such that $[\mathbb{Q}_{p}(t_{k}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Since $(s_{n},\delta_{n})$ is a minimal pair, we have $v_{p}(s_{n}-t_{k})<\delta_{n}$, so that
$\lambda_{k}=\overline{v}_{E}(X-t_{k})=\overline{v}_{E}(X-s_{n}+s_{n}-t_{k})<\overline{v}_{E}(X-s_{n})=\delta_{n}$ (2.6)
Conversely, let $n\in\mathbb{N}$. We need to show that there exists $k\in\mathbb{N}$ such that $\delta_{n}\leq\lambda_{k}$. For all $m$ sufficiently large we have
$\overline{v}_{E}(X-s_{n})=v_{p}(t_{m}-s_{n})=\overline{v}_{E}(t_{m}-X+X-s_{n})$
and since $n$ is fixed and $\overline{v}_{E}(t_{m}-X)=\lambda_{m}$ is strictly increasing, it follows that $\overline{v}_{E}(t_{m}-X)=\lambda_{m}>\overline{v}_{E}(X-s_{n})$ for all such $m$’s.
Hence, $\text{Br}(E)=\text{Br}(E^{\prime})$.
Finally, we need to show that, if $k\in\mathbb{N}$, then there exist $n_{0},m_{0}\in\mathbb{N}$ such that for each $n\geq n_{0}$ and $m\geq m_{0}$, we have $v_{p}(t_{n}-s_{m})>\lambda_{k}$. Let $n_{0}$ be the smallest integer such that $[\mathbb{Q}_{p}(t_{k}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n_{0}}):\mathbb{Q}_{p}]$. As in (2.6) above, $\lambda_{k}<v_{E}(X-s_{n_{0}})=\delta_{n_{0}}$. Let now $m>k$ and $n\geq n_{0}$. Then,
$v(t_{m}-s_{n})=\overline{v}_{E}(t_{m}-X+X-s_{n})>\lambda_{k}$
since $\overline{v}_{E}(t_{m}-X)=\lambda_{m}>\lambda_{k}$ and $\overline{v}_{E}(X-s_{n})\geq\overline{v}_{E}(X-s_{n_{0}})=\delta_{n_{0}}>\lambda_{k}$. Hence, $E$ and $E^{\prime}$ are equivalent.
By [23, Proposition 5.3], $\overline{\mathbb{Z}_{p}}_{E}=\overline{\mathbb{Z}_{p}}_{E^{\prime}}$, so, in particular, $\mathbb{Z}_{p,E}=\mathbb{Z}_{p,E^{\prime}}$. The final claim follows by Proposition 2.4. ∎
The following proposition describes the value group and the residue field of a residually algebraic torsion extension $W$ of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$. By Theorem 2.5, $W$ is equal to $\mathbb{Z}_{p,E}$, for some strongly stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$. We keep the notation of Remark 2.2.
Proposition 2.7.
Let $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ be a stacked sequence and $W=\mathbb{Z}_{p,E}$. Then we have
$\bigcup_{n\in\mathbb{N}}\Gamma_{n}=\Gamma_{w},\;\;\bigcup_{n\in\mathbb{N}}k_{n}=k_{w}.$
Proof.
Let $w=v_{E}$ be the valuation associated to $\mathbb{Z}_{p,E}$ and $\overline{v}_{E}$ the valuation associated to $\overline{\mathbb{Z}_{p}}_{E}$.
Since $E$ is of transcendental type, for each $f\in\mathbb{Q}_{p}[X]$, $v_{E}(f)=v(f(s_{n}))$ for all $n$ sufficiently large (see §1.1). It follows that for each $\phi\in\mathbb{Q}_{p}(X)$, $\phi=f/g$, for some $f,g\in\mathbb{Q}_{p}[X]$, $v_{E}(\phi)=v_{E}(f)-v_{E}(g)$ is in $\Gamma_{n}$ for all $n$ sufficiently large. Hence, $\Gamma_{w}\subseteq\bigcup_{n}\Gamma_{n}$. Conversely, let $n\in\mathbb{N}$ and $f\in\mathbb{Q}_{p}[X]$ be of degree smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Then, each root $\alpha_{i}$ of $f(X)$ in $\overline{\mathbb{Q}_{p}}$ has degree smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and so, since $(s_{n},\delta_{n})$ is a minimal pair, we have
$v_{p}(s_{n}-\alpha_{i})<\delta_{n}$ (2.8)
which implies that
$\overline{v}_{E}(X-\alpha_{i})=\overline{v}_{E}(X-s_{n}+s_{n}-\alpha_{i})=v_{p}(s_{n}-\alpha_{i})$ (2.9)
and so
$v_{E}(f(X))=\sum_{i}\overline{v}_{E}(X-\alpha_{i})=\sum_{i}v_{p}(s_{n}-\alpha_{i})=v_{p}(f(s_{n}))$ (2.10)
which shows that $\Gamma_{n}\subseteq\Gamma_{w}$. Note that $\overline{v}_{E}(X-\alpha_{i})=v(s_{m}-\alpha_{i})$ for each $m\geq n$ and so $v_{E}(f(X))=v(f(s_{m}))$ for each $m\geq n$.
Let now $n\in\mathbb{N}$ and $\overline{c}=\overline{f(s_{n})}\in k_{n}$, for some $f(s_{n})\in O_{n}^{*}$ where $f\in\mathbb{Q}_{p}[X]$ has degree strictly smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. In particular, $\overline{c}\not=0$. As in (2.10), $v_{E}(f(X))=v(f(s_{m}))=0$ for each $m\geq n$. Let $\alpha_{i}$ be a root of $f(X)$ in $\overline{\mathbb{Q}_{p}}$. Then by (2.9), $\overline{v}_{E}(X-\alpha_{i})=v_{p}(s_{n}-\alpha_{i})=v_{p}(d_{i})$ for some $d_{i}\in\overline{\mathbb{Q}_{p}}$. Then
$\overline{v}_{E}\left(\frac{(X-\alpha_{i})/d_{i}}{(s_{n}-\alpha_{i})/d_{i}}-1\right)=\overline{v}_{E}\left(\frac{X-s_{n}}{s_{n}-\alpha_{i}}\right)=\delta_{n}-v_{p}(s_{n}-\alpha_{i})>0$
where the last inequality holds by (2.8). Therefore, $(X-\alpha_{i})/d_{i}$ and $(s_{n}-\alpha_{i})/d_{i}$ coincide over the residue field of $W$. In particular,
$\frac{f(X)}{f(s_{n})}=\prod_{i}\frac{(X-\alpha_{i})}{(s_{n}-\alpha_{i})}=\prod_{i}\frac{(X-\alpha_{i})/d_{i}}{(s_{n}-\alpha_{i})/d_{i}}$ (2.11)
and since each factor of the last product has residue $\overline{1}$ in $W$, it follows that $f(X)$ and $f(s_{n})$ coincide over the residue field of $W$ (which contains both $f(X)$ and $f(s_{n})$). Since $f\in\mathbb{Z}_{p,E}=W$, this shows that $k_{n}$ is contained in the residue field $k_{w}$ of $W$.
Conversely, let $\phi=f/g\in W\subset\mathbb{Q}_{p}(X)$, for some $f,g\in\mathbb{Q}_{p}[X]$. Let $\alpha_{i},\beta_{j}$ be the roots in $\overline{\mathbb{Q}_{p}}$ of $f$ and $g$, respectively. There exists $n\in\mathbb{N}$ such that $[\mathbb{Q}_{p}(\alpha_{i}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and $[\mathbb{Q}_{p}(\beta_{j}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ for all $i$ and $j$. Hence, as in (2.9) we have
$\overline{v}_{E}(X-\alpha_{i})=v_{p}(s_{n}-\alpha_{i}),\;\;\overline{v}_{E}(X-\beta_{j})=v(s_{n}-\beta_{j}),\forall i,j$
which again as in (2.10) shows that
$v_{E}(\phi(X))=v(\phi(s_{n}))$
Moreover, this last equation holds if we replace $s_{n}$ by $s_{m}$, for all $m\geq n$. If $v_{E}(\phi(X))=0$, then as in (2.11) one can show that $\phi(X)$ and $\phi(s_{n})$ coincide over the residue field of $W$, so that $k_{w}\subseteq k_{n}$. ∎
The following corollary gives a further characterization of the residue field and the value group of a residually algebraic torsion extension $W$ of $\mathbb{Z}_{p}$: either the residue field of $W$ is an infinite algebraic extension of $\mathbb{F}_{p}$, or the value group $\Gamma_{w}$ is non-discrete.
Corollary 2.12.
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ and let $e=e(W|\mathbb{Z}_{p})$ and $f=f(W|\mathbb{Z}_{p})$ be the ramification index and the residue field degree, respectively. Then $e\cdot f=\infty$.
Proof.
By Theorem 2.5, there exists a stacked sequence $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{p,E}$. By Proposition 2.7, $\Gamma_{w}=\bigcup_{n}\Gamma_{n}$ and $k_{w}=\bigcup_{n}k_{n}$. Remark 2.2 shows that either the sequence $\{e_{n}=[\Gamma_{n}:\mathbb{Z}]\}_{n\in\mathbb{N}}$ or $\{f_{n}=[k_{n}:\mathbb{F}_{p}]\}_{n\in\mathbb{N}}$ is unbounded, therefore, either $e=e(W|\mathbb{Z}_{p})$ or $f=f(W|\mathbb{Z}_{p})$ is infinite. ∎
The following proposition is analogous to [5, Proposition 2.3]. It shows that the sequence of ramification indexes, residue field degrees and gauges attached to a residually algebraic torsion extension $W$ of $\mathbb{Z}_{p}$ do not depend on the strongly stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{p,E}$ (Theorem 2.5).
Proposition 2.13.
Let $W\subset\mathbb{Q}_{p}(X)$ be a residually algebraic torsion extension of $\mathbb{Z}_{p}$. Let $E=\{s_{n}\}_{n\in\mathbb{N}},E^{\prime}=\{t_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ be strongly stacked sequences with gauges $\{\delta_{n}\}_{n\in\mathbb{N}},\{\delta^{\prime}_{n}\}_{n\in\mathbb{N}}$, respectively, such that $W=\mathbb{Z}_{p,E}=\mathbb{Z}_{p,E^{\prime}}$. Then, for each $n\in\mathbb{N}$ we have:
  • i)
    $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$ and $\delta_{n}=\delta_{n}^{\prime}$.
  • ii)
    $e_{s_{n}}=e_{t_{n}}$ and $f_{s_{n}}=f_{t_{n}}$.
Proof.
Without loss of generality, we may assume that $\overline{W}=\overline{\mathbb{Z}_{p}}_{E}=\overline{\mathbb{Z}_{p}}_{E^{\prime}}$.
Let $w$ be a valuation associated to $\overline{W}$.
We prove i). We have $s_{0},t_{0}\in\mathbb{Q}_{p}$. There exists $n\in\mathbb{N}$, $n\geq 1$, such that $w(X-s_{n-1})\leq w(X-t_{0})<w(X-s_{n})$, otherwise $t_{0}$ would be a pseudo-limit of $E$, which is not possible. In particular, $v_{p}(s_{n}-t_{0})=w(s_{n}-X+X-t_{0})=w(X-t_{0})\geq w(X-s_{n-1})=\delta_{n-1}$. If $n>1$ we have $[\mathbb{Q}_{p}(t_{0}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ so by iii) of Definition 2.1 we have that $v_{p}(s_{n}-t_{0})=v_{p}(s_{n-1}-t_{0})<v_{p}(s_{n}-s_{n-1})=\delta_{n-1}$ which is impossible. Hence, $n=1$, so $v_{p}(s_{1}-t_{0})=w(X-t_{0})\geq w(X-s_{0})$. Reversing the roles of $s_{0},t_{0}$, we get the other inequality, so $w(X-s_{0})=w(X-t_{0})=\delta_{0}=\delta_{0}^{\prime}$.
Let $n\in\mathbb{N}$ and suppose that for each $m\leq n$ we have $[\mathbb{Q}_{p}(s_{m}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]$ and $\delta_{m}=\delta_{m}^{\prime}$.
Since $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$, by ii) of Definition 2.1 we have $v_{p}(t_{n+1}-s_{n})\leq v_{p}(t_{n+1}-t_{n})=\delta_{n}^{\prime}=\delta_{n}$. Now,
$v_{p}(t_{n+1}-s_{n})=v_{p}(t_{n+1}-t_{n}+t_{n}-s_{n})\geq\delta_{n}$
since $v_{p}(t_{n}-s_{n})=w(t_{n}-X+X-s_{n})\geq\delta_{n}=\delta_{n}^{\prime}$. This implies that $v_{p}(t_{n+1}-s_{n})=\delta_{n}$ and so $(s_{n},t_{n+1})$ is distinguished. Moreover, we have
$v_{p}(t_{n+1}-s_{n+1})=w(t_{n+1}-X+X-s_{n+1})>\delta_{n}=\delta_{n}^{\prime}=v_{p}(t_{n+1}-s_{n}).$
Now, if $[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$, then since $(s_{n},t_{n+1})$ is distinguished, we would have $v_{p}(s_{n+1}-t_{n+1})\leq v_{p}(t_{n+1}-s_{n})$, which is impossible. Hence, $[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]\geq[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$. The other inequality is proved in a symmetrical way, so $[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$.
Suppose now that $w(X-s_{n+1})<w(X-t_{n+1})$. Then $v_{p}(s_{n+2}-t_{n+1})=w(s_{n+2}-X+X-t_{n+1})>w(X-s_{n+1})=v_{p}(s_{n+2}-s_{n+1})$ which is not possible since $(s_{n+1},s_{n+2})$ is distinguished. Hence, $w(X-s_{n+1})\geq w(X-t_{n+1})$. The other inequality is proved similarly, so $\delta_{n+1}=\delta_{n+1}^{\prime}$ as claimed.
We prove now ii). For each $n\in\mathbb{N}$, let $\Gamma_{n},\Gamma^{\prime}_{n}$ and $k_{n},k_{n}^{\prime}$ be the value groups and residue fields, respectively, of $\mathbb{Q}_{p}(s_{n})$ and $\mathbb{Q}_{p}(t_{n})$. Let $e_{n}=e_{s_{n}}$, $e_{n}^{\prime}=e_{t_{n}}$, $f_{n}=f_{s_{n}}$, $f^{\prime}_{n}=f_{t_{n}}$.
Clearly, $e_{0}=e_{0}^{\prime}$ and $f_{0}=f_{0}^{\prime}$, since $s_{0},t_{0}\in\mathbb{Q}_{p}$.
Let $n\geq 1$. If $f\in\mathbb{Q}_{p}[X]$ has degree strictly smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$, then by (2.10) we have $w(f(X))=v_{p}(f(s_{n}))$ and also $w(f(X))=v_{p}(f(t_{n}))$, so $v_{p}(f(s_{n}))=v_{p}(f(t_{n}))$. This proves that $\Gamma_{n}=\Gamma_{n}^{\prime}$ and so $e_{n}=e_{n}^{\prime}$.
Suppose now that $f\in\mathbb{Q}_{p}[X]$ of degree strictly smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$ is such that $v_{p}(f(s_{n}))=v_{p}(f(t_{n}))=0$. In particular, $w(f(X))=0$ by (2.10). By (2.11) and the analogous equation where $s_{n}$ is replaced by $t_{n}$, we get that $f(s_{n}),f(t_{n})$ have the same residue as $f(X)$, so in particular, $k_{n}=k_{n}^{\prime}$. Therefore, $f_{n}=f_{n}^{\prime}$. ∎
2.2 Residually algebraic extensions of $\mathbb{Z}_{p}$ which are DVRs
In this section we characterize DVRs of $\mathbb{Q}_{p}(X)$ extending $\mathbb{Z}_{p}$ such that the residue field extension is algebraic, necessarily of infinite degree by Corollary 2.12; this fact has already been noted in a different way in [21, p. 4217]. We will see in §2.3 that there is no such restriction on the residue field degree for DVRs of $\mathbb{Q}(X)$ which are residually algebraic extensions of $\mathbb{Z}_{(p)}$ (see Corollary 2.28).
Given $\alpha\in\mathbb{C}_{p}$, we denote by $\mathbb{O}_{p,\alpha}$ the unique valuation domain of $\mathbb{Q}_{p}(\alpha)$ lying over $\mathbb{Z}_{p}$ (i.e., $\mathbb{O}_{p,\alpha}=\mathbb{O}_{p}\cap\mathbb{Q}_{p}(\alpha)$). We also set
$\mathbb{Z}_{p,\alpha}=\{\phi\in\mathbb{Q}_{p}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\},$
which is a valuation domain of $\mathbb{Q}_{p}(X)$ and coincides with the previous definition if $\alpha\in\overline{\mathbb{Q}_{p}}$.
Proposition 2.14.
Let $\alpha\in\mathbb{C}_{p}$ be a transcendental element. Then there exists a Cauchy stacked sequence $E\subseteq\overline{\mathbb{Q}_{p}}$ converging to $\alpha$. Moreover, the valued fields $(\mathbb{Q}_{p}(X),\mathbb{Z}_{p,\alpha})$ and $(\mathbb{Q}_{p}(\alpha),\mathbb{O}_{p,\alpha})$ are isomorphic. In particular, the ramification index $e(\mathbb{Z}_{p,\alpha}|\mathbb{Z}_{p})$ is equal to $e_{\alpha}$, the residue field degree $f(\mathbb{Z}_{p,\alpha}|\mathbb{Z}_{p})$ is equal to $f_{\alpha}$ and $e_{\alpha}\cdot f_{\alpha}=\infty$.
Note that the last condition implies that either $e_{\alpha}$ or $f_{\alpha}$ is infinite. It can happen that exactly one of these two quantities is finite (see Theorem 2.21).
Proof.
The proof of the first claim follows also by [5, Proposition 2.2], but we give here a different proof based on the previous results.
By Theorem 2.5, there exists a stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$ such that $\mathbb{Z}_{p,\alpha}=\mathbb{Z}_{p,E}$. Since the valuation domains $\overline{\mathbb{Z}_{p}}_{E},\overline{\mathbb{Z}_{p}}_{\alpha}\subset\overline{\mathbb{Q}_{p}}(X)$ contracts down to $\mathbb{Q}_{p}(X)$ to the same valuation domain, there exists $\sigma\in\text{Gal}(\overline{\mathbb{Q}_{p}}/\mathbb{Q}_{p})$ such that $\sigma(\overline{\mathbb{Z}_{p}}_{\alpha})=\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}=\overline{\mathbb{Z}_{p}}_{E}$. By [23, Proposition 5.3], $E$ is then a Cauchy sequence converging to $\sigma(\alpha)$. Since $\mathbb{Z}_{p,\alpha}=\mathbb{Z}_{p,\sigma(\alpha)}$, without loss of generality we may assume that $E$ converges to $\alpha$.
Since $\alpha$ is transcendental over $\mathbb{Q}_{p}$, the evaluation homomorphism $\text{ev}_{\alpha}:\mathbb{Q}_{p}(X)\to\mathbb{Q}_{p}(\alpha)$, $\phi(X)\mapsto\phi(\alpha)$, is an isomorphism. It is easy to see that $\text{ev}_{\alpha}(\mathbb{Z}_{p,\alpha})=\mathbb{O}_{p,\alpha}$. Hence, $\mathbb{Z}_{p,\alpha}$ and $\mathbb{O}_{p,\alpha}$ have the same ramification indexes and residue field degrees over $\mathbb{Z}_{p}$.
Finally, the last claim follows by Corollary 2.12. ∎
Remark 2.15.
By Proposition 2.14, we may conclude that in general,
for $\text{for }\alpha\in\mathbb{C}_{p},\text{ we have }e_{\alpha}\cdot f_{\alpha}<\infty\text{ if and only if }\alpha\in\overline{\mathbb{Q}_{p}}.$
The next lemma may be well-known, but lacking a reference we give a short proof.
Lemma 2.16.
Let $p\in\mathbb{Z}$ be a prime, $K_{1},K_{2}$ finite extensions of $\mathbb{Q}_{p}$ and $L=K_{1}K_{2}$ the compositum. Let $e_{1}$ be the ramification index of $K_{1}$ over $\mathbb{Q}_{p}$ and $e$ the ramification index of $L$ over $K_{2}$. Then $e\leq e_{1}$.
Proof.
If $K_{1}$ is a tame extension of $\mathbb{Q}_{p}$, then the ramification index of $L$ over $\mathbb{Q}_{p}$ is equal to lcm$\{e(K_{1}|\mathbb{Q}_{p}),e(K_{2}|\mathbb{Q}_{p})\}$ (see for example [7]), so $e$ divides $e_{1}$ and the claim is true.
We give a self-contained proof which works in general. Let $L^{\prime}$ be the normal closure of $L$ over $\mathbb{Q}_{p}$ and $I$ the inertia group of the maximal ideal $M_{L^{\prime}}$ of $O_{L^{\prime}}$ over $\mathbb{Z}_{p}$. Let $G_{i}$ be the Galois group $\text{Gal}(L^{\prime}|K_{i})$, for $i=1,2$ and $G$ the Galois group $\text{Gal}(L^{\prime}|L)$. Since $L=K_{1}K_{2}$, we have $G=G_{1}\cap G_{2}$. The inertia group of $M_{L^{\prime}}$ over $M_{K_{1}}$ is equal to $I\cap G_{1}$ and the inertia group of $M_{L^{\prime}}$ over $M_{L}$ is equal to $I\cap G$. We have
$e=\frac{e(L^{\prime}|K_{2})}{e(L^{\prime}|L)}=\frac{\#(I\cap G_{2})}{\#(I\cap G)},\;\;e_{1}=\frac{e(L^{\prime}|\mathbb{Q}_{p})}{e(L^{\prime}|K_{1})}=\frac{\#I}{\#(I\cap G_{1})}$
Note that $I\cap G=(I\cap G_{1})\cap(I\cap G_{2})$. Therefore, the claim follows by the following general fact for finite groups: given a finite group $G$ with two subgroups $H_{1},H_{2}$, we have
#$\frac{\#H_{2}}{\#(H_{1}\cap H_{2})}=[H_{2}:H_{1}\cap H_{2}]\leq\frac{\#G}{\#H_{1}}=[G:H_{1}]$
which follows immediately, since the map $h_{2}H_{1}\cap H_{2}\mapsto h_{2}H_{1}$ from the set $\{h_{2}(H_{1}\cap H_{2})\mid h_{2}\in H_{2}\}$ of left cosets of $H_{1}\cap H_{2}$ in $H_{2}$ to the set $\{gH_{1}\mid g\in G\}$ of left cosets of $H_{1}$ in $G$ is injective. ∎
The following result is analogous to [21, Theorem 2.5].
Theorem 2.17.
Let $W$ be a DVR of $\mathbb{Q}_{p}(X)$ which is a residually algebraic extension of $\mathbb{Z}_{p}$. Then there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $W=\mathbb{Z}_{p,\alpha}$.
Proof.
Note that, by Corollary 2.12, the residue field of $W$ is an infinite algebraic extension of $\mathbb{F}_{p}$.
By Theorem 2.5, there exists a stacked sequence $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{p,E}$. By assumption, the ramification index $e(W|\mathbb{Z}_{p})=e$ is finite. By Remark 2.2 and Proposition 2.7, there exists $n_{0}\in\mathbb{N}$ such that $\Gamma_{w}=\Gamma_{n}=\Gamma_{n_{0}}$ for each $n\geq n_{0}$. Equivalently, $e_{n}=e_{n_{0}}=e$ for each $n\geq n_{0}$. Let $n\geq n_{0}$. Note that $\delta_{n}=v_{p}(s_{n+1}-s_{n})\in\Gamma_{O_{K_{n}}}$, where $K_{n}=\mathbb{Q}_{p}(s_{n},s_{n+1})$. Note that the ramification index of $\mathbb{Q}_{p}(s_{i})$ over $\mathbb{Q}_{p}$ is equal to $e$, for $i=n,n+1$. By Lemma 2.16, the ramification index of $K_{n}$ over $\mathbb{Q}_{p}$ is bounded by $e^{2}$. If $d=\prod_{i=1}^{e^{2}}i$, then $d\delta_{n}\in\mathbb{Z}$, for each $n\geq n_{0}$. This shows that the gauge $\{\delta_{n}\}_{n\in\mathbb{N}}$ of $E$ has bounded denominator, so $\delta_{n}\nearrow\infty$, thus $E$ is Cauchy and converges to a (unique) element $\alpha$ of $\mathbb{C}_{p}\setminus\overline{\mathbb{Q}_{p}}$, since $E$ is of transcendental type by Lemma 2.3. In particular, $W=\mathbb{Z}_{p,\alpha}$. ∎
Remark 2.18.
We say that an element $\alpha\in\mathbb{C}_{p}$ has bounded ramification if the extension $\mathbb{Q}_{p}(\alpha)\supseteq\mathbb{Q}_{p}$ has finite ramification. We denote by $\mathbb{C}_{p}^{\text{br}}$ the set of all elements of $\mathbb{C}_{p}$ of bounded ramification; clearly, $\overline{\mathbb{Q}_{p}}\subset\mathbb{C}_{p}^{\text{br}}$. A transcendental element $\alpha\in\mathbb{C}_{p}$ has bounded ramification if and only if the set of ramification indexes $\{e_{n}\}_{n\in\mathbb{N}}$ attached to a stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$ converging to $\alpha$ is bounded; in fact, by Theorem 2.17, the integer $e$ such that $e=e_{n}$ for all $n$ sufficiently large is equal to $e(\mathbb{Q}_{p}(\alpha)|\mathbb{Q}_{p})$.
We remark that not all the transcendental elements $\alpha\in\mathbb{C}_{p}$ have bounded ramification. For example, according to [12] there exist generic transcendental elements $t\in\mathbb{C}_{p}$ for $\mathbb{C}_{p}$, that is, the completion of $\mathbb{Q}_{p}(t)$ is equal to $\mathbb{C}_{p}$. In particular, the value group of the unique valuation of $\mathbb{O}_{p,t}$ is equal to $\mathbb{Q}$, so the corresponding ramification index is $\infty$. Hence, by Proposition 2.7, $\mathbb{Z}_{p,t}$ has value group equal to $\mathbb{Q}$ and therefore the set of ramification indexes $\{e_{n}\}_{n\in\mathbb{N}}$ is unbounded.
We show in Theorem 2.21 that given any algebraic extension $k$ of $\mathbb{F}_{p}$ and group $\Gamma$ such that $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$, there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$, provided that either $[k:\mathbb{F}_{p}]$ is infinite or $\Gamma$ is not discrete (this condition being necessary by Corollary 2.12).
Lemma 2.19.
Let $l$ be an infinite algebraic extension of $\mathbb{Q}_{p}$ such that $e(l|\mathbb{Q}_{p})$ is finite. Then $l$ is contained in the maximal unramified extension $K^{\text{unr}}$ of a finite extension $K$ of $\mathbb{Q}_{p}$.
Proof.
For each $n\in\mathbb{N}$, let $\mathbb{Q}_{p}^{(n)}$ be the compositum of all the extensions of $\mathbb{Q}_{p}$ of degree bounded by $n$. Clearly, $\overline{\mathbb{Q}_{p}}=\bigcup_{n\in\mathbb{N}}\mathbb{Q}_{p}^{(n)}$ and $\mathbb{Q}_{p}^{(n)}\subset\mathbb{Q}_{p}^{(n+1)}$ for each $n\in\mathbb{N}$. Since $\mathbb{Q}_{p}$ has only finitely many extensions of bounded degree, $\mathbb{Q}_{p}^{(n)}=\mathbb{Q}_{p}(t_{n})$ for some $t_{n}\in\overline{\mathbb{Q}_{p}}$. Now, for each $n\in\mathbb{N}$, we let $\mathbb{Q}_{p}(t_{n})\cap l=\mathbb{Q}_{p}(s_{n})$ for some $s_{n}\in l$. Clearly, $l=\bigcup_{n\in\mathbb{N}}\mathbb{Q}_{p}(s_{n})$ and $\mathbb{Q}_{p}(s_{n})\subset\mathbb{Q}_{p}(s_{n+1})$, for each $n\in\mathbb{N}$. Since $\Gamma_{s_{n}}\subseteq\Gamma_{s_{n+1}}\subseteq\Gamma_{l}$ for each $n\in\mathbb{N}$ and $\Gamma_{l}$ is discrete by assumption, there exists $n_{0}\in\mathbb{N}$ such that $\Gamma_{s_{n}}=\Gamma_{s_{n_{0}}}$ for each $n\geq n_{0}$. Therefore, if $K=\mathbb{Q}_{p}(s_{n_{0}})$, then $s_{n}\in K^{\text{unr}}$ for each $n\geq n_{0}$, so that $l\subseteq K^{\text{unr}}$. ∎
The next lemma shows that a transcendental element $t$ of $\mathbb{C}_{p}$ with bounded ramification arise as the limit of sequences contained in the maximal unramified extension $K^{\text{unr}}$ of a finite extension $K$ of $\mathbb{Q}_{p}$. We don’t know whether there exists a stacked sequence in $K^{\text{unr}}$ which converges to $t$.
Proposition 2.20.
Let $t\in\mathbb{C}_{p}^{\text{br}}$. Then $t$ is the limit of a sequence contained in the maximal unramified extension of a finite extension of $\mathbb{Q}_{p}$.
Proof.
By [12, Theorem 1], the completion of $\widehat{\mathbb{Q}_{p}(t)}\cap\overline{\mathbb{Q}_{p}}$ is equal to $\widehat{\mathbb{Q}_{p}(t)}$. In particular, there exists a Cauchy sequence $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\widehat{\mathbb{Q}_{p}(t)}\cap\overline{\mathbb{Q}_{p}}$ converging to $t$. Now, since $\mathbb{Q}_{p}(t)\subset\widehat{\mathbb{Q}_{p}(t)}$ and $\widehat{\mathbb{Q}_{p}(t)}\cap\overline{\mathbb{Q}_{p}}\subset\widehat{\mathbb{Q}_{p}(t)}$ are immediate extensions, it follows that $\widehat{\mathbb{Q}_{p}(t)}\cap\overline{\mathbb{Q}_{p}}$ has value group $\Gamma_{t}$ and residue field $k_{t}$. By Lemma 2.19, $\widehat{\mathbb{Q}_{p}(t)}\cap\overline{\mathbb{Q}_{p}}$ is contained in the maximal unramified extension of a finite extension of $\mathbb{Q}_{p}$. The statement follows. ∎
The following result is not new, see for example [15, Lemma 2], the present proof is different because it employs the notion of stacked sequence.
Theorem 2.21.
Let $k$ be an algebraic extension of $\mathbb{F}_{p}$ and $\Gamma$ a totally ordered group with $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$, such that either $[k:\mathbb{F}_{p}]$ or $[\Gamma:\mathbb{Z}]$ is infinite (the last condition is equivalent to $\Gamma$ being not discrete). Then there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $k_{\alpha}=k$ and $\Gamma_{\alpha}=\Gamma$. In particular, $\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$.
Note that, by Corollary 2.12, the last claim shows that $[k:\mathbb{F}_{p}]\cdot[\Gamma:\mathbb{Z}]=\infty$ is necessary.
Proof.
Since $\overline{\mathbb{F}_{p}}$ is countable, we may suppose that $k=\bigcup_{n\in\mathbb{N}}k_{n}$, where $k_{n}$ is a finite extension of $\mathbb{F}_{p}$, $k_{n}\subseteq k_{n+1}$ and $k_{0}=\mathbb{F}_{p}$. Similarly, $\Gamma=\bigcup_{n\in\mathbb{N}}\Gamma_{n}$, where $\Gamma_{n}$ is a discrete group, $\Gamma_{n}\subseteq\Gamma_{n+1}$ and $\Gamma_{0}=\mathbb{Z}$. Let $f=[k:\mathbb{F}_{p}]$ and $e=[\Gamma:\mathbb{Z}]$; then, either $e$ or $f$ is infinite. Without loss of generality, we may assume that for each $n$, $[k_{n+1}:k_{n}][\Gamma_{n+1}:\Gamma_{n}]>1$.
For each $n\in\mathbb{N}$, there exists a local field $K_{n}=\mathbb{Q}_{p}(s_{n})$ with residue field $k_{n}$ and value group $\Gamma_{n}$. By induction, we may also assume that $K_{n}\subset K_{n+1}$. Let $\{\lambda_{n}\}_{n\in\mathbb{N}}\subset\mathbb{Q}$ be a strictly increasing sequence in $\mathbb{Q}$ which is unbounded and $\lambda_{0}<\delta_{0}=v(s_{1}-s_{0})$.
We define now a sequence $E=\{t_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ such that for each $n\in\mathbb{N}$, $n\geq 1$, we have
  • i)
    $\mathbb{Q}_{p}(t_{n})=\mathbb{Q}_{p}(s_{n})$;
  • ii)
    $(t_{n-1},\delta_{n-1}=v_{p}(t_{n}-t_{n-1}))$ is a minimal pair;
  • iii)
    $\delta_{n-1}>\lambda_{n-1}$.
In particular, $E$ is a stacked sequence by conditions i) and ii) and Cauchy by condition iii) and the assumption on $\{\lambda_{n}\}_{n\in\mathbb{N}}$.
We set $t_{0}=s_{0}\in\mathbb{Q}_{p}$, $t_{1}=s_{1}\not\in\mathbb{Q}_{p}$ and $\delta_{0}=v_{p}(t_{1}-t_{0})$. Note that $(t_{0},\delta_{0})$ is a minimal pair. We proceed by induction on $n$. We assume that for all $m<n$ we have chosen $t_{m}\in\overline{\mathbb{Q}_{p}}$ such that conditions i), ii), iii) above are satisfied.
We now show how to choose $t_{n}$. We choose $a_{n}\in\mathbb{Q}_{p}$, $a_{n}\not=0$, such that
$v_{p}(a_{n})>\max\{\omega(t_{n-1})-v_{p}(s_{n}),\lambda_{n-1}-v_{p}(s_{n})\}$
We then set
$t_{n}=a_{n}s_{n}+t_{n-1}.$
Note that $\mathbb{Q}_{p}(t_{n})\subseteq\mathbb{Q}_{p}(s_{n})$, since by induction $\mathbb{Q}_{p}(t_{n-1})=\mathbb{Q}_{p}(s_{n-1})$ and the last field is contained in $\mathbb{Q}_{p}(s_{n})$. Now, since $\delta_{n-1}=v_{p}(t_{n}-t_{n-1})>\omega(t_{n-1})$, it follows by Krasner’s Lemma that $\mathbb{Q}_{p}(t_{n-1})\subseteq\mathbb{Q}_{p}(t_{n})$. This containment and the fact that $s_{n}=\frac{t_{n}-t_{n-1}}{a_{n}}$ show that $s_{n}$ is in $\mathbb{Q}_{p}(t_{n})$, so that $\mathbb{Q}_{p}(t_{n})=\mathbb{Q}_{p}(s_{n})$. Moreover, note also that $\delta_{n-1}>\lambda_{n-1}$. Hence, $E=\{t_{n}\}_{n\in\mathbb{N}}$ is a stacked sequence which is Cauchy, so $E$ converges to a transcendental element $\alpha$ of $\mathbb{C}_{p}$. By Proposition 2.7, $\mathbb{Z}_{p,E}=\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$, as wanted. By Proposition 2.14, $\mathbb{Z}_{p,\alpha}$ is isomorphic to $\mathbb{O}_{p,\alpha}$, so it follows the $\Gamma_{\alpha}=\Gamma$ and $k_{\alpha}=k$. ∎
Remark 2.22.
We remark that without condition iii) above in the proof of Theorem 2.21, in general we may only conclude that there exists a stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$ (which may not be Cauchy) such that the valuation domain $\mathbb{Z}_{p,E}$ has residue field $k$ and value group $\Gamma$. If instead $\Gamma$ is discrete by assumption, condition iii) is not necessary: in fact, there exists $n_{0}\in\mathbb{N}$ such that $\Gamma_{n}=\Gamma_{n_{0}}=\Gamma$ for all $n\geq n_{0}$; that is, $K_{n}=\mathbb{Q}_{p}(s_{n})$ is an unramified extension of $K_{n_{0}}$ for all $n>n_{0}$. Hence, $E\subset\bigcup_{n\in\mathbb{N}}K_{n}$ is Cauchy, and so $\mathbb{Z}_{p,E}=\mathbb{Z}_{p,\alpha}$, where $\alpha\in\mathbb{C}_{p}^{\text{br}}$ is the transcendental limit of $E$.
We close this section showing that the statement of [12, Proposition 1] is wrong, namely, in general the completion of $\mathbb{Q}_{p}(X)$ with respect to a residually algebraic torsion extension $W$ of $\mathbb{Z}_{p}$ may not be a subfield of $\mathbb{C}_{p}$. The mistake is due to the fact that if $W=\mathbb{Z}_{p,E}$ for some pseudo-convergent sequence $E\subset\overline{\mathbb{Q}_{p}}$ of transcendental type, then $X$ is a pseudo-limit of $E$ with respect to $w$ and may not be a limit (that is, $E$ may not be Cauchy).
Proposition 2.23.
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$. Then the completion $\widehat{\mathbb{Q}_{p}(X)}$ with respect to $W$ is a subfield of $\mathbb{C}_{p}$ if and only if there exists a transcendental element $\alpha$ in $\mathbb{C}_{p}$ such that $W=\mathbb{Z}_{p,\alpha}$.
Proof.
By Theorem 2.5, there exists a pseudo-convergent sequence $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ of transcendental type such that $W=\mathbb{Z}_{p,E}$.
Suppose that $\widehat{\mathbb{Q}_{p}(X)}\subseteq\mathbb{C}_{p}$. In particular, $X\in\mathbb{C}_{p}$ and so there exists a Cauchy sequence $F=\{t_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ which tends to $X$. Since $\mathbb{Q}_{p}(X)\subset\overline{\mathbb{Q}_{p}}(X)$ is an algebraic extension and $\mathbb{C}_{p}$ is algebraically closed, then also the completion of $\overline{\mathbb{Q}_{p}}(X)$ with respect to $\overline{W}=\overline{\mathbb{Z}_{p}}_{E}$ is contained in $\mathbb{C}_{p}$. Without loss of generality, we may suppose that the restriction of $v_{p}$ to $\overline{\mathbb{Q}_{p}}(X)$ is equal to $\overline{w}$. In particular, $\overline{w}(X-t_{n})=v_{p}(X-t_{n})\nearrow\infty$. Since $E$ is of transcendental type, for each $n$ there exists $m_{0}$ such that $\overline{w}(X-t_{n})<\overline{w}(X-s_{m})$ for each $m\geq m_{0}$. This shows that the gauge of $E$ tends to infinity, thus $E$ is Cauchy; in particular, $E$ converges to a transcendental element $\alpha\in\mathbb{C}_{p}$. Therefore, $W=\mathbb{Z}_{p,\alpha}$.
Conversely, let $W=\mathbb{Z}_{p,\alpha}$ for some transcendental element $\alpha\in\mathbb{C}_{p}$. Then, by Proposition 2.14, the completion $\widehat{\mathbb{Q}_{p}(X)}$ with respect to $\mathbb{Z}_{p,\alpha}$ is isomorphic to the completion of $\mathbb{Q}_{p}(\alpha)$ and therefore can be identified to a subfield of $\mathbb{C}_{p}$. ∎
In particular, if $W=\mathbb{Z}_{p,E}$ for some stacked non-Cauchy sequence $E\subset\overline{\mathbb{Q}_{p}}$, then $\widehat{\mathbb{Q}_{p}(X)}$ is not contained in $\mathbb{C}_{p}$.
2.3 Residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$
We now characterize residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$. We remark that such a valuation domain may have an extension to $\mathbb{Q}_{p}(X)$ which is a residually algebraic extension of $\mathbb{Z}_{p}$ but is not torsion. For example, let $\alpha\in\overline{\mathbb{Q}_{p}}$ be transcendental over $\mathbb{Q}$, then $\mathbb{Z}_{(p),\alpha}$ is torsion but $\mathbb{Z}_{p,\alpha}$ is not (the one dimensional valuation overring of $\mathbb{Z}_{p,\alpha}$ is $\mathbb{Q}_{p}[X]_{(p_{\alpha}(X))}$, where $p_{\alpha}(X)$ is the minimal polynomial of $\alpha$ over $\mathbb{Q}_{p}$).
Given $\alpha\in\mathbb{C}_{p}$, we consider the following valuation domain of $\mathbb{Q}(X)$:
$\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$
which is just the contraction to $\mathbb{Q}(X)$ of $\mathbb{Z}_{p,\alpha}$ considered in §2.1. Similarly, if $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ is a pseudo-convergent sequence of transcendental type, then we set
$\mathbb{Z}_{(p),E}=\{\phi\in\mathbb{Q}(X)\mid\phi(s_{n})\in\overline{\mathbb{Z}_{p}},\text{ for all sufficiently large }n\in\mathbb{N}\}$
which is equal to $\mathbb{Z}_{p,E}\cap\mathbb{Q}(X)$.
The next proposition is analogous to Proposition 2.4, and characterizes residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ in terms of pseudo-convergent sequences of $\overline{\mathbb{Q}_{p}}$ which are of transcendental type over $\mathbb{Q}$; clearly, every pseudo-convergent sequence of transcendental type of $\overline{\mathbb{Q}_{p}}$ belongs to this class. As a particular case, we find again part of the result of [21, Theorem 2.5].
Proposition 2.24.
Let $p\in\mathbb{P}$ and let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$. Then there exists a pseudo-convergent sequence $E\subset\overline{\mathbb{Q}_{p}}$ of transcendental type over $\mathbb{Q}$ such that $W=\mathbb{Z}_{(p),E}$. More precisely, let $e,f$ be the ramification index and residue field degree of $W$ over $\mathbb{Z}_{(p)}$, respectively. Let $\widehat{\mathbb{Q}(X)}$ be the completion of $\mathbb{Q}(X)$ with respect to the $W$-adic topology. Then the following conditions are equivalent:
  1. 1.
    $\widehat{\mathbb{Q}(X)}$ is a finite extension of $\mathbb{Q}_{p}$.
  2. 2.
    $X$ is algebraic over $\mathbb{Q}_{p}$.
  3. 3.
    $W=\mathbb{Z}_{(p),\alpha}$, for some $\alpha\in\overline{\mathbb{Q}_{p}}$ transcendental over $\mathbb{Q}$.
  4. 4.
    $ef<\infty$.
If either one of these conditions holds, then the sequence $E$ above is Cauchy and converges to $\alpha$ (and $E$ is therefore of algebraic type over $\mathbb{Q}_{p}$). Moreover, we have $\Gamma_{w}=\Gamma_{\alpha}$ and $k_{w}=k_{\alpha}$.
If $ef=\infty$, then $E\subset\overline{\mathbb{Q}_{p}}$ is of transcendental type over $\mathbb{Q}_{p}$ and $\mathbb{Z}_{(p),E}\subset\mathbb{Z}_{p,E}$ is an immediate extension.
Proof.
Note that since $W$ is a torsion extension of $\mathbb{Z}_{(p)}$, the $p$-adic completion $\mathbb{Q}_{p}$ of $\mathbb{Q}$ is contained in $\widehat{\mathbb{Q}(X)}$ (see for example the arguments given in the proof of [2, Corollary 2.6]).
If $\widehat{\mathbb{Q}(X)}$ is a finite extension of $\mathbb{Q}_{p}$ then clearly $X$ is algebraic over $\mathbb{Q}_{p}$, so 1. implies 2. If $X$ is algebraic over $\mathbb{Q}_{p}$, we may identify $X$ with some $\alpha\in\overline{\mathbb{Q}_{p}}$; $\mathbb{Q}_{p}(\alpha)$ is a finite extension of $\mathbb{Q}_{p}$, hence complete. So, $\widehat{\mathbb{Q}(X)}=\mathbb{Q}_{p}(\alpha)$. As in the proof of [21, Theorem 2.5] it follows easily that $W=\mathbb{Z}_{(p),\alpha}$. Therefore, 2. implies 3.
If $W=\mathbb{Z}_{(p),\alpha}$ for some $\alpha\in\overline{\mathbb{Q}_{p}}$ transcendental over $\mathbb{Q}$, then by [21, Proposition 2.2], $ef<\infty$, so 3. implies 4. Finally, 4. implies 3. by [21, Lemma 2.4].
Note that if $E\subset\overline{\mathbb{Q}_{p}}$ is a pseudo-convergent sequence converging to $\alpha$ (hence, $E$ is a Cauchy sequence, which is of transcendental type over $\mathbb{Q}$ but of algebraic type over $\mathbb{Q}_{p}$), then $W=\mathbb{Z}_{(p),E}=\mathbb{Z}_{(p),\alpha}$ (see §1.1).
The claims about the value group and residue field of $\mathbb{Z}_{(p),\alpha}$ follow by [21, Proposition 2.2].
If $ef=\infty$ then $X$ is transcendental over $\mathbb{Q}_{p}$ by the previous part of the proof; in particular, the field of rational functions $\mathbb{Q}_{p}(X)$ is contained in $\widehat{\mathbb{Q}(X)}$. If $\widetilde{W}=\widehat{W}\cap\mathbb{Q}_{p}(X)$, then $\widetilde{W}$ is a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$, so by Theorem 2.5 there exists a stacked sequence $E\subset\overline{\mathbb{Q}_{p}}$ such that $\widetilde{W}=\mathbb{Z}_{p,E}$ (by Lemma 2.3, $E$ is a pseudo-convergent sequence of transcendental type, necessarily unbounded). Restricting down to $\mathbb{Q}(X)$ we get $W=\mathbb{Z}_{(p),E}$. Finally, since $W\subset\widehat{W}$ is an immediate extension, it follows that $\mathbb{Z}_{(p),E}\subset\mathbb{Z}_{p,E}$ is an immediate extension, too. Hence, the value group and residue field of $\mathbb{Z}_{(p),E}$ are the same as those of $\mathbb{Z}_{p,E}$, respectively (see Proposition 2.7). ∎
The following statement is the analogous of Proposition 2.23 for residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$.
Corollary 2.25.
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$. Then the completion $\widehat{\mathbb{Q}(X)}$ with respect to $W$ is a subfield of $\mathbb{C}_{p}$ if and only if there exists $\alpha\in\mathbb{C}_{p}$, transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}$.
Proof.
According to Proposition 2.24, when passing to the completion, either $X$ is algebraic over $\mathbb{Q}_{p}$ or $X$ is transcendental over $\mathbb{Q}_{p}$ and consequently either $\widehat{\mathbb{Q}(X)}\subset\overline{\mathbb{Q}_{p}}\subset\mathbb{C}_{p}$ or $\mathbb{Q}_{p}(X)\subset\widehat{\mathbb{Q}(X)}$. In the first case, $W=\mathbb{Z}_{(p),\alpha}$ for some $\alpha\in\overline{\mathbb{Q}_{p}}\subset\mathbb{C}_{p}$ transcendental over $\mathbb{Q}$. In the second case, $\widehat{\mathbb{Q}_{p}(X)}=\widehat{\mathbb{Q}(X)}$, where the completion of $\mathbb{Q}_{p}(X)$ is considered with respect to the valuation domain $\widetilde{W}=\widehat{W}\cap\mathbb{Q}_{p}(X)$. In particular, by Proposition 2.23, we get that $\widehat{\mathbb{Q}(X)}\subseteq\mathbb{C}_{p}$ if and only if there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $W=\mathbb{Z}_{(p),\alpha}$. ∎
In particular, if $W=\mathbb{Z}_{(p),E}$ for some stacked non-Cauchy sequence $E\subset\overline{\mathbb{Q}_{p}}$, then $\widehat{\mathbb{Q}(X)}$ is not contained in $\mathbb{C}_{p}$.
The following result is the analogous of Theorem 2.21 for building residually algebraic torsion extensions $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ with prescribed residue field $k$ and value group $\Gamma$. Note that, contrary to that Theorem, now we are not assuming anymore that $[k:\mathbb{F}_{p}]\cdot[\Gamma:\mathbb{Z}]=\infty$.
Theorem 2.26.
Let $k$ be an algebraic extension of $\mathbb{F}_{p}$ and $\Gamma$ a totally ordered group such that $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$. Then there exists $\alpha\in\mathbb{C}_{p}$, transcendental over $\mathbb{Q}$, such that $\mathbb{Z}_{(p),\alpha}$ has residue field $k$ and value group $\Gamma$.
Proof.
Let $e=[\Gamma:\mathbb{Z}]$ and $f=[k:\mathbb{F}_{p}]$. If $ef<\infty$, then it is well known that there exists $\alpha\in\overline{\mathbb{Q}_{p}}$ transcendental over $\mathbb{Q}$ such that $\mathbb{O}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$. Hence, by [21, Proposition 2.2], $\mathbb{Z}_{(p),\alpha}$ is the desired extension of $\mathbb{Z}_{(p)}$.
If $ef=\infty$, then, by Theorem 2.21, there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$. Clearly, $\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$ is a residually algebraic torsion extension of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$. Moreover, by Proposition 2.14, $\mathbb{Z}_{p,\alpha}=\mathbb{Z}_{p,E}$ for some stacked Cauchy sequence $E\subset\overline{\mathbb{Q}_{p}}$ which converges to $\alpha$. In particular, $\mathbb{Z}_{(p),\alpha}=\mathbb{Z}_{(p),E}$. By the last part of Proposition 2.24, $\mathbb{Z}_{(p),E}\subset\mathbb{Z}_{p,E}$ is an immediate extension, so $\mathbb{Z}_{(p),\alpha}$ has residue field $k$ and value group $\Gamma$. ∎
Now we are able to describe the DVRs of $\mathbb{Q}(X)$ which are residually algebraic extensions of $\mathbb{Z}_{(p)}$, for some $p\in\mathbb{P}$. We recall that every $\sigma\in G_{\mathbb{Q}_{p}}=\text{Gal}(\overline{\mathbb{Q}_{p}}/\mathbb{Q}_{p})$ extends uniquely to a continuous automorphism of $\mathbb{C}_{p}$, see [5, §3]. Given $\alpha,\beta\in\mathbb{C}_{p}$, we say that $\alpha,\beta$ are conjugate (over $\mathbb{Q}_{p}$) if there exists $\sigma\in G_{\mathbb{Q}_{p}}=\text{Gal}(\overline{\mathbb{Q}_{p}}/\mathbb{Q}_{p})$ such that $\sigma(\alpha)=\beta$; the orbit of an element $\alpha\in\mathbb{C}_{p}$ is finite if and only if $\alpha\in\overline{\mathbb{Q}_{p}}$ (see [5, Remark 3.2]).
We prove first the following lemma.
Lemma 2.27.
Let $p\in\mathbb{P}$ and $W$ a valuation domain of $\mathbb{Q}_{p}(X)$ such that $W\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$ for some $\alpha\in\mathbb{C}_{p}$. Then $W=\mathbb{Z}_{p,\alpha}$.
Proof.
Let $n\geq 0$ be an integer such that $p^{n}\cdot\alpha=\alpha_{0}\in\mathbb{O}_{p}$. The field isomorphism $X\mapsto\frac{X}{p^{n}}$ maps $\mathbb{Z}_{(p),\alpha}$ to $\mathbb{Z}_{(p),\alpha_{0}}$ and $\mathbb{Z}_{p,\alpha}$ to $\mathbb{Z}_{p,\alpha_{0}}$, respectively. Hence, in order to prove the statement, without loss of generality, we may assume that $\alpha\in\mathbb{O}_{p}$.
Let $w$ be a valuation associated to $W$. We note first that since $X\in\mathbb{Z}_{(p),\alpha}$, it follows that $w(X)\geq 0$. Let $f\in W\cap\mathbb{Q}_{p}[X]$, say $f(X)=\sum_{i=0}^{d}\alpha_{i}X^{i}$. Then for $g(X)=\sum_{i=0}^{d}a_{i}X^{i}\in\mathbb{Q}[X]$ we have
$w(f-g)\geq\min_{0\leq i\leq d}\{v_{p}(\alpha_{i}-a_{i})+iw(X)\}.$
Therefore, if we choose $a_{i}\in\mathbb{Q}$ sufficiently $v_{p}$-adically close to $\alpha_{i}$ for each $i=0,\ldots,d$, we have $w(f-g)\geq 0$. In particular, $g\in W\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$. The polynomial $h=f-g$ is in $\mathbb{Z}_{p}[X]$; therefore $f(\alpha)=h(\alpha)+g(\alpha)\in\mathbb{O}_{p}$, so that $f\in\mathbb{Z}_{p,\alpha}$. Therefore $W\cap\mathbb{Q}_{p}[X]\subseteq\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}_{p}[X]$. Similarly, one can easily show that the other containment holds, so $W\cap\mathbb{Q}_{p}[X]=\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}_{p}[X]$. In the same way, $M_{W}\cap\mathbb{Q}_{p}[X]=M_{p,\alpha}\cap\mathbb{Q}_{p}[X]$, where $M_{W}$ and $M_{p,\alpha}$ are the maximal ideals of $W$ and $\mathbb{Z}_{p,\alpha}$, respectively.
Let now $\psi\in\mathbb{Z}_{p,\alpha}$; since $\mathbb{Z}_{p}[X]\subset\mathbb{Q}_{p}[X]\cap\mathbb{Z}_{p,\alpha}$, we may suppose that $\psi=\frac{f}{g}$, where $f,g\in\mathbb{Q}_{p}[X]\cap\mathbb{Z}_{p,\alpha}$. Clearly, $g(\alpha)\not=0$; then, there exists $n\in\mathbb{N}$, $n\geq 1$ and $c\in\mathbb{Q}_{p}$, $c\not=0$, such that $v_{p}(c)+v_{p}(g(\alpha)^{n})=0$. We consider then the rational function $\psi^{n}=\frac{cf^{n}}{cg^{n}}=\frac{f_{1}}{g_{1}}$, which still is in $\mathbb{Z}_{p,\alpha}$. Note that $f_{1}\in\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}_{p}[X]=W\cap\mathbb{Q}_{p}[X]$ and $g_{1}\in\mathbb{Z}_{p,\alpha}^{*}\cap\mathbb{Q}_{p}[X]=W^{*}\cap\mathbb{Q}_{p}[X]$, since $v_{p,\alpha}(f_{1})\geq v_{p,\alpha}(g_{1})=0$ (the denotes the set of units of the valuation domains). In particular,
$w(f_{1})\geq 0=w(g_{1})$
which proves that $\psi^{n}\in W$. Since $W$ is integrally closed, it follows that $\psi\in W$. Hence, $\mathbb{Z}_{p,\alpha}\subseteq W$. The equality follows because both rings are extensions of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ and in the case $\alpha$ is algebraic over $\mathbb{Q}_{p}$, the one-dimensional valuation overring of $\mathbb{Z}_{p,\alpha}$ is non-unitary (i.e., $\mathbb{Q}_{p}[X]_{(q)}$, where $q\in\mathbb{Q}_{p}[X]$ is the minimal polynomial of $\alpha$). ∎
Corollary 2.28.
Let $W$ be a DVR of $\mathbb{Q}(X)$ which is a residually algebraic extension of $\mathbb{Z}_{(p)}$ for some $p\in\mathbb{P}$. Then there exists $\alpha\in\mathbb{C}_{p}^{\text{br}}$, transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}$. The element $\alpha$ belongs to $\overline{\mathbb{Q}_{p}}$ if and only if the residue field extension $\mathbb{Z}/p\mathbb{Z}\subseteq W/M$ is finite.
Moreover, for $\alpha,\beta\in\mathbb{C}_{p}$, we have $\mathbb{Z}_{(p),\alpha}=\mathbb{Z}_{(p),\beta}$ if and only if there exists $\sigma\in G_{\mathbb{Q}_{p}}$ such that $\sigma(\alpha)=\beta$.
Proof.
Let $f=[W/M:\mathbb{Z}/p\mathbb{Z}]$. If $f<\infty$, then the claim follows by [21, Theorem 2.5] and corresponds to the first case of Proposition 2.24: $W=\mathbb{Z}_{(p),\alpha}$, for some $\alpha\in\overline{\mathbb{Q}_{p}}$ which is transcendental over $\mathbb{Q}$. If $f=\infty$, then we are in the last case of Proposition 2.24, so $W=\mathbb{Z}_{(p),E}$, for some pseudo-convergent sequence in $\overline{\mathbb{Q}_{p}}$ of transcendental type. As in the proof of Proposition 2.24, we denote by $\widehat{W}$ the completion of $W$; since the ramification index $e(W\mid\mathbb{Z}_{(p)})$ is finite, $\widetilde{W}=\widehat{W}\cap\mathbb{Q}_{p}(X)$ is a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ which is a DVR, so by Theorem 2.17, $\widetilde{W}=\mathbb{Z}_{p,\alpha}$ for some $\alpha\in\mathbb{C}_{p}^{\text{br}}\setminus\overline{\mathbb{Q}_{p}}$. Hence, $W=\widetilde{W}\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$. Note that $\alpha$ is transcendental over $\mathbb{Q}_{p}$, hence also over $\mathbb{Q}$.
We prove now the final claim. Suppose there exists $\sigma\in G_{\mathbb{Q}_{p}}$ such that $\sigma(\alpha)=\beta$. If $\phi\in\mathbb{Z}_{(p),\alpha}$, then $\phi(\alpha)$ is defined and belongs to $\mathbb{O}_{p}$. In particular, $\sigma(\phi(\alpha))=\phi(\sigma(\alpha))=\phi(\beta)\in\overline{\mathbb{Z}_{p}}$. Hence, $\mathbb{Z}_{(p),\alpha}\subseteq\mathbb{Z}_{(p),\beta}$ and the other containment is proved in a symmetrical way.
Conversely, suppose that $\mathbb{Z}_{(p),\alpha}=\mathbb{Z}_{(p),\beta}$. By Lemma 2.27, it follows that $\mathbb{Z}_{p,\alpha}=\mathbb{Z}_{p,\beta}$. Note that the last two valuation domains are the contraction to $\mathbb{Q}_{p}(X)$ of the valuation domains $\overline{\mathbb{Z}_{p}}_{\alpha}=\{\phi\in\overline{\mathbb{Q}_{p}}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\},\overline{\mathbb{Z}_{p}}_{\beta}=\{\phi\in\overline{\mathbb{Q}_{p}}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$ of $\overline{\mathbb{Q}_{p}}(X)$, respectively. By [6, 4, Chapt. VI, §8, 6., Corollary 1], there exists a $\mathbb{Q}_{p}(X)$-automorphism $\sigma$ of $\overline{\mathbb{Q}_{p}}(X)$ such that $\sigma(\overline{\mathbb{Z}_{p}}_{\alpha})=\overline{\mathbb{Z}_{p}}_{\beta}$. It is easy to check that $\sigma(\overline{\mathbb{Z}_{p}}_{\alpha})=\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}$. In particular, $\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}=\overline{\mathbb{Z}_{p}}_{\beta}$. If $\sigma(\alpha)-\beta\not=0$, let $c\in\overline{\mathbb{Z}_{p}}$ be such that $v_{p}(c)>v_{p}(\sigma(\alpha)-\beta)$. Let $a\in\overline{\mathbb{Q}_{p}}$ be such that $v_{p}(a-\sigma(\alpha))\geq v_{p}(c)$. Then the polynomial $\frac{X-a}{c}$ is in $\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}$ and not in $\overline{\mathbb{Z}_{p}}_{\beta}$, a contradiction. ∎
Note that, for a DVR $W$ as in the statement of Corollary 2.28, there exists $\alpha\in\mathbb{O}_{p}\subset\mathbb{C}_{p}$ of bounded ramification such that $W=\mathbb{Z}_{(p),\alpha}$ if and only if $X\in W$. This last condition occurs for example if $W$ is an overring of $\mathbb{Z}[X]$.
3 Polynomial Dedekind domains
In order to describe the family of Dedekind domains lying between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$, we briefly recall the notion of integer-valued polynomials on algebras (see [8, 24], for example). Let $D$ be an integral domain with quotient field $K$ and $A$ a torsion-free $D$ algebra. We embed $K$ and $A$ into the extended $K$-algebra $B=A\otimes_{D}K$, and this allows us to evaluate polynomials over $K$ at elements of $A$. If $f\in K[X]$ and $a\in A$ are such that $f(a)\in A$, then we say that $f$ is integer-valued at $a$. In general, given a subset $S$ of $A$, we denote by
$\textnormal{Int}_{K}(S,A)=\{f\in K[X]\mid f(s)\in A,\forall s\in S\}$
the ring of integer-valued polynomials over $S$. We omit the subscript $K$ if $A=D$.
In our setting, let $\mathcal{O}=\prod_{p\in\mathbb{P}}\mathbb{O}_{p}\subset\prod_{p\in\mathbb{P}}\mathbb{C}_{p}$. Given $\alpha=(\alpha_{p})\in\prod_{p\in\mathbb{P}}\mathbb{C}_{p}$ and $f\in\mathbb{Q}[X]$, then $f(\alpha)=(f(\alpha_{p}))$, which is an element of $\prod_{p\in\mathbb{P}}\mathbb{C}_{p}$. If $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}$ is a subset of $\prod_{p}\mathbb{C}_{p}$, then
$\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})=\{f\in\mathbb{Q}[X]\mid f(\alpha)\in\mathcal{O},\forall\alpha\in\underline{E}\}$
that is, a polynomial $f$ is in $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ if $f(\alpha_{p})\in\mathbb{O}_{p}$ for each $\alpha_{p}\in E_{p}$ and $p\in\mathbb{P}$. By an argument similar to [8, Remark 6.3] there is no loss in generality to suppose that a subset of $\prod_{p\in\mathbb{P}}\mathbb{C}_{p}$ is of the form $\prod_{p\in\mathbb{P}}E_{p}$, when dealing with such rings of integer-valued polynomials.
We remark that we have the following representation for the ring $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ as an intersection of valuation overrings (see [19, (2.2)], for example):
$\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})=\bigcap_{p\in\mathbb{P}}\bigcap_{\alpha_{p}\in E_{p}}\mathbb{Z}_{(p),\alpha_{p}}\cap\bigcap_{q\in\mathcal{P}^{\text{irr}}}\mathbb{Q}[X]_{(q)}$ (3.1)
where $\mathcal{P}^{\text{irr}}$ denotes the set of irreducible polynomials in $\mathbb{Q}[X]$. By [21, Proposition 2.2], the valuation domain $\mathbb{Z}_{(p),\alpha_{p}}$ of $\mathbb{Q}(X)$ has rank one if and only if $\alpha_{p}$ is transcendental over $\mathbb{Q}$, and has rank $2$ otherwise (in the last case, note that necessarily $\alpha\in\overline{\mathbb{Q}_{p}}$).
A totally similar argument to [19, Lemma 2.5] shows that, for $p\in\mathbb{P}$, we have
$(\mathbb{Z}\setminus p\mathbb{Z})^{-1}(\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O}))=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})=\{f\in\mathbb{Q}[X]\mid f(E_{p})\subseteq\mathbb{O}_{p}\}$
We also need to recall the following definition introduced in [19].
Definition 3.2.
We say that a subset $\underline{E}$ of $\mathcal{O}$ is polynomially factorizable if, for each $g\in\mathbb{Z}[X]$ and $\alpha=(\alpha_{p})\in\underline{E}$, there exist $n,d\in\mathbb{Z}$, $n,d\geq 1$ such that $\frac{g(\alpha)^{n}}{d}$ is a unit of $\mathcal{O}$, that is, $v_{p}(\frac{g(\alpha_{p})^{n}}{d})=0$, $\forall p\in\mathbb{P}$.
The next theorem characterizes which rings of integer-valued polynomials $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ are Dedekind domains. Given $p\in\mathbb{P}$ and a subset $E_{p}$ of $\mathbb{C}_{p}$, we say that $E_{p}$ has finitely many $G_{\mathbb{Q}_{p}}=\text{Gal}(\overline{\mathbb{Q}_{p}}/\mathbb{Q}_{p})$-orbits if $E_{p}$ contains finitely many equivalence classes under the relation of conjugacy over $\mathbb{Q}_{p}$ (we stress that $E_{p}$ may not necessarily contain a full $G_{\mathbb{Q}_{p}}$-orbit). By Corollary 2.28, this condition holds if and only if the set $\{\mathbb{Z}_{(p),\alpha_{p}}\mid\alpha_{p}\in E_{p}\}$ is finite. Furthermore, if $E_{p}\subseteq\overline{\mathbb{Q}_{p}}$, then the number of $G_{\mathbb{Q}_{p}}$-orbits is finite if and only if $E_{p}$ is a finite set.
Theorem 3.3.
Let $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}\subset\prod_{p}\mathbb{C}_{p}$. Then $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ is a Dedekind domain if and only if, for each prime $p$, $E_{p}$ is a subset of $\mathbb{C}_{p}^{\text{br}}$ of transcendental elements over $\mathbb{Q}$ with finitely many $G_{\mathbb{Q}_{p}}$-orbits and $\underline{E}$ is polynomially factorizable.
Moreover, if the above conditions hold, then the class group of $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ is isomorphic to the direct sum of the class groups $\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$, $p\in\mathbb{P}$, and if $E_{p}=\{\alpha_{1},\ldots,\alpha_{n}\}$ where the $\alpha_{i}$’s are pairwise non-conjugate over $\mathbb{Q}_{p}$, then $\text{Cl}(\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p}))=\mathbb{Z}/e\mathbb{Z}\oplus\mathbb{Z}^{n-1}$, where $e$ is the greatest common divisors of the ramifications indexes of $\alpha_{i}$ over $\mathbb{Q}_{p}$.
In particular, assuming that $E_{p}$ is formed by pairwise non-conjugate elements over $\mathbb{Q}_{p}$ for each $p\in\mathbb{P}$, $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ is a PID if and only if, $\underline{E}$ is polynomially factorizable and for each $p\in\mathbb{P}$, $E_{p}$ contains at most one element $\alpha_{p}\in\mathbb{O}_{p}\cap\mathbb{C}_{p}^{\text{br}}$, such that $\alpha_{p}$ is transcendental over $\mathbb{Q}$ and unramified over $\mathbb{Q}_{p}$.
Proof.
Let $R=\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$.
Suppose that the above conditions on $\underline{E}$ are satisfied. By (3.1), $R$ is equal to an intersection of DVRs. Moreover, $R$ has finite character, that is, for every non-zero $f\in R$, $f$ belongs to finitely many maximal ideals of the family of DVRs appearing in (3.1): in fact, if $f(X)=\frac{g(X)}{n}$, for some $g\in\mathbb{Z}[X]$ and $n\in\mathbb{Z}$, $n\not=0$, then $f$ is divisible only by finitely many $q\in\mathcal{P}^{\text{irr}}$; since $\underline{E}$ is polynomially factorizable, by [19, Lemma 2.12], the set $\{p\in\mathbb{P}\mid\exists\alpha_{p}\in E_{p},v_{p}(g(\alpha_{p}))>0\}$ is finite, so that $f$ belongs to finitely many maximal ideals of the family $\mathbb{Z}_{(p),\alpha_{p}}$, $\alpha_{p}\in E_{p},p\in\mathbb{P}$. Hence, $R$ is a Krull domain.
Suppose that $R$ is not a Dedekind domain. By [11, Proposition 2.2], there exists a maximal ideal $M\subset R$ of height strictly bigger than one. If $M\cap\mathbb{Z}=(0)$, then, since $\mathbb{Z}[X]\subseteq R\subseteq\mathbb{Q}[X]$, it follows that $R_{\mathbb{Z}\setminus\{0\}}=\mathbb{Q}[X]$ and $12\leq htM=ht(M_{\mathbb{Z}\setminus\{0\}})\leq dim(\mathbb{Q}[X])=$ 1, a contradiction. Hence, $M\cap\mathbb{Z}=p\mathbb{Z}$, for some $p\in\mathbb{P}$. If we now localize at $p$, we have that $(\mathbb{Z}\setminus p\mathbb{Z})^{-1}R=R_{p}=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$ which is a Dedekind domain by [9, Theorem]. So $(\mathbb{Z}\setminus p\mathbb{Z})^{-1}M\subset R_{p}$ cannot have dimension strictly bigger than one, a contradiction.
Conversely, suppose that $R$ is a Dedekind domain. In particular, for each $p\in\mathbb{P}$, $(\mathbb{Z}\setminus p\mathbb{Z})^{-1}R=R_{p}=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$ is a Dedekind domain, so $\{\mathbb{Z}_{(p),\alpha_{p}}\mid\alpha_{p}\in E_{p}\}$ is a finite set of DVRs (because $p$ is contained in only finitely many maximal ideals of these valuation overrings) which implies that $E_{p}$ is a subset of $\mathbb{C}_{p}^{\text{br}}$ of transcendental elements over $\mathbb{Q}$ and $E_{p}$ has finitely many $G_{\mathbb{Q}_{p}}$-orbits. Since every polynomial of $R$ is contained in only finitely many maximal ideals, it follows easily that $\underline{E}$ is polynomially factorizable.
Finally, suppose that $R$ is a Dedekind domain. As in [19, Lemma 2.14], we have $\text{Cl}(R)=\bigoplus_{p\in\mathbb{P}}\text{Cl}(R_{p})$, where $R_{p}=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$ for $p\in\mathbb{P}$. The claim about the class group of $R_{p}$ follows by [19, Proposition 2.10] or by [9, Theorem], since, for each $p\in\mathbb{P}$, we are assuming that $E_{p}=\{\alpha_{1},\ldots,\alpha_{n}\}$ is formed by pairwise non-conjugate elements over $\mathbb{Q}_{p}$.
The claim about when $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathbb{O})$ is a PID is now straightforward. ∎
Let ^¯$\overline{\widehat{\mathbb{Z}}}=\prod_{p\in\mathbb{P}}\overline{\mathbb{Z}_{p}}$. In [19, Theorem 2.17] we show that if $R$ is a Dedekind domain between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ such that the residue fields of prime characteristic are finite fields, then $R=\textnormal{Int}_{\mathbb{Q}}(\underline{E},\overline{\widehat{\mathbb{Z}}})$, for some $\underline{E}=\prod_{p}E_{p}\subset\overline{\widehat{\mathbb{Z}}}$ such that $\underline{E}$ is polynomially factorizable and for each $p\in\mathbb{P}$, $E_{p}$ is a finite subset of $\overline{\mathbb{Z}_{p}}$ of transcendental elements over $\mathbb{Q}$. Now, we are able to complete the classification of the Dedekind domains $R$, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, without any restriction on the residue fields.
Theorem 3.4.
Let $R$ be a Dedekind domain such that $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$. Then $R$ is equal to $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$, for some polynomially factorizable subset $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}\subset\mathcal{O}$, such that, for each prime $p$, $E_{p}\subset\mathbb{O}_{p}\cap\mathbb{C}_{p}^{\text{br}}$ is a finite set of transcendental elements over $\mathbb{Q}$.
Proof.
Note first that, by [20, Theorem 3.14], no valuation overring of $W$ of $R$ can be a residually transcendental extension of $W\cap\mathbb{Q}$, since for such a valuation domain $W$, the domain $W\cap\mathbb{Q}[X]$ is not Prüfer. Hence, for each prime ideal $P\subset R$ such that $P\cap\mathbb{Z}=p\mathbb{Z}$, $p\in\mathbb{P}$, $R_{P}$ is a DVR of $\mathbb{Q}(X)$ which is a residually algebraic extension of $\mathbb{Z}_{(p)}$. By Corollary 2.28, there exists $\alpha\in\mathbb{O}_{p}\cap\mathbb{C}_{p}^{\text{br}}$ such that $R_{p}=\mathbb{Z}_{(p),\alpha}$. Let $E_{p}$ be the subset of $\mathbb{C}_{p}^{\text{br}}$ formed by all such $\alpha_{p}$’s. Note that since $p$ is contained in only finitely many maximal ideals $P$ of $R$, it follows that $E_{p}$ is a finite set; moreover, each element of $E_{p}$ is transcendental over $\mathbb{Q}$, since $R_{P}$ is a DVR. It now follows that
$R=\bigcap_{p\in\mathbb{P}}\bigcap_{\begin{subarray}{c}P\subset R\\ P\cap\mathbb{Z}=p\mathbb{Z}\end{subarray}}R_{P}\cap\mathbb{Q}[X]=\bigcap_{p\in\mathbb{P}}\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})=\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O}).$
The rest of the statement follows by Theorem 3.3. ∎
Finally, the next corollary describes the PIDs among the family of Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$.
Corollary 3.5.
Let $R$ be a PID such that $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$. Then $R$ is equal to $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$, for some $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}\subset\mathcal{O}$, such that, for each prime $p$, $E_{p}$ contains at most one element $\alpha_{p}\in\mathbb{O}_{p}\cap\mathbb{C}_{p}^{\text{br}}$, such that $\alpha_{p}$ is transcendental over $\mathbb{Q}$ and unramified over $\mathbb{Q}_{p}$ and $\underline{E}=\{\alpha=(\alpha_{p})\}$ is polynomially factorizable.
Proof.
By Theorem 3.4, the ring $R$ is equal to $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$, for some polynomially factorizable subset $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}\subset\mathcal{O}$, such that, for each prime $p$, $E_{p}\subset\mathbb{O}_{p}\cap\mathbb{C}_{p}^{\text{br}}$ is a set of transcendental elements over $\mathbb{Q}$ with finitely many $G_{\mathbb{Q}_{p}}$-orbits. Since by hypothesis the class group of $R$ is trivial, it follows by Theorem 3.3 that for each $p\in\mathbb{P}$, $E_{p}$ contains at most one element, which is transcendental over $\mathbb{Q}$ and unramified over $\mathbb{Q}_{p}$. ∎
Remark 3.6.
As we mentioned in the Introduction, given a group $G$ which is the direct sum $G$ of a countable family of finitely generated abelian groups, there exists a Dedekind domain $R$ between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ with class group $G$ ([19, Theorem 3.1]). The domain $R$ of that construction is equal to $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ for some polynomially factorizable subset $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}$, where $E_{p}$ is a finite subset of $\overline{\mathbb{Q}_{p}}$ of transcendental elements over $\mathbb{Q}$. In particular, $R$ has finite residue fields of prime characteristic ([19, Theorem 2.17]); the reason is that the valuation overrings $\mathbb{Z}_{(p),\alpha_{p}}$ of $R$ in (3.1) have finite residue fields precisely because $\alpha_{p}$ is chosen in $\overline{\mathbb{Q}_{p}}$ for each $p\in\mathbb{P}$ (Proposition 2.24).
Now, by means of Theorem 2.26, with the same method used in [19, Theorem 3.1], we can build a Dedekind domain $R$, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, with prescribed class group $G$ as above and prescribed residue fields of prime characteristic, which can be finite or infinite algebraic extensions of the prime field $\mathbb{F}_{p}$, according to whether the the above elements $\alpha_{p}\in E_{p}\subset\mathbb{C}_{p}^{\text{br}}$ transcendental over $\mathbb{Q}$, are either algebraic or transcendental over $\mathbb{Q}_{p}$.
Acknowledgments
The author wishes to thank the anonymous referee for his/her remarks.